### v1

# Adaptive Testing for LLM Evaluation: A Psychometric Alternative to Static Benchmarks

Abstract
Large language model evaluation requires thousands of benchmark items, making evaluations expensive and slow. Existing methods compute average accuracy across fixed item sets, treating all items equally despite varying quality and informativeness. We present ATLAS  an adaptive testing framework using Item Response Theory (IRT) to estimate model ability through Fisher information-guided item selection. Our analysis of five major benchmarks reveals that 3-6% of items exhibit negative discrimination, indicating annotation errors that corrupt static evaluation. ATLAS achieves 90% item reduction while maintaining measurement precision: on HellaSwag (5,608 items), we match full-benchmark estimates using only 42 items with 0.154 MAE. Our framework maintains item exposure rates below 10% and test overlap at 16-27%, compared to static benchmarks where every model sees all items (100% exposure). Among 4,000+ tested models, IRT ranks differ from accuracy ranks: models with the same accuracy get different IRT scores, and 23-31% of all models shift by more than 10 rank positions. Code and calibrated item banks are available at https://github.com/Peiyu-Georgia-Li/ATLAS.git.
**footnotetext: These authors contributed equally.
1 Introduction
Large language model (LLM) evaluation relies on benchmarks with tens of thousands of test items. These benchmarks impose high computational costs and evaluation cycles spanning days or weeks. Despite benchmark growth—some proposals exceed 100,000 items—evaluation practice remains static: models receive average accuracy scores across fixed item sets. This approach ignores statistical information in large datasets and raises questions about efficiency and validity.
Current evaluation faces three fundamental limitations. First, average scores hide meaningful differences between models with distinct error patterns, particularly for lower-performing models where small ability differences are obscured by measurement noise. Second, static benchmarks create vulnerability to data contamination as items leak into pretraining corpora, enabling high scores through memorization rather than genuine capability. Third, evaluating complete benchmarks is inefficient, treating poorly discriminative items as equally informative as high-quality questions.
To address these limitations, we propose ATLAS (Adaptive Testing for LLM Ability Scoring), an adaptive evaluation framework based on computerized adaptive testing (CAT; (lord1980, ; wainer2000computerized, ; weiss1982improving, )). ATLAS first calibrates benchmark items using three-parameter logistic (3PL) IRT models to estimate item difficulty, discrimination, and guessing parameters (birnbaum1968some, ; hambleton1991fundamentals, ). Then, rather than administering fixed item sets, ATLAS dynamically selects items with maximum Fisher information for each model’s current estimated ability, terminating when precision thresholds are reached. This approach directly addresses all three limitations: Fisher information-guided selection provides precise ability estimates that distinguish models with identical accuracy, dynamic item selection reduces contamination through diverse test forms, and adaptive termination achieves reliable evaluation with significantly fewer items.
We validate our framework across five major benchmarks (WinoGrande, TruthfulQA, HellaSwag, GSM8K, ARC). ATLAS achieves competitive or superior performance to static baselines: on TruthfulQA, it attains the lowest MAE (0.067) using only 51 items; on HellaSwag, it matches MetaBench accuracy (0.154 MAE) while requiring 3× fewer items (39 vs 93). Across benchmarks, ATLAS requires 30-78 items on average compared to thousands in full evaluation, while IRT-based estimates reveal systematic rank reordering that exposes limitations of accuracy-based evaluation.
Our contributions are: (1) We demonstrate fundamental limitations of average-score evaluation and establish the benefits of psychometric ability estimates for LLM evaluation. (2) We develop ATLAS, the first large-scale adaptive testing framework for LLMs, achieving 90% item reduction while maintaining measurement precision and natural contamination resistance through $10\%$ item exposure rates. (3) We provide comprehensive psychometric analysis of five major benchmarks, revealing that IRT-based estimates cause systematic rank reordering ($23-31\%$ of models shift more than 10 positions) and identifying that $3-6\%$ of items exhibit negative discrimination due to annotation errors, with calibrated item banks released for reproducible evaluation.
2 Related Work
2.1 Background
Large-scale benchmarks such as WinoGrande, TruthfulQA, HellaSwag, GSM8K and ARC have become standard for LLM evaluation but impose significant computational costs. These static approaches suffer from benchmark saturation, data contamination, and item redundancy, raising concerns about their ability to reflect genuine model capabilities.
Item Response Theory (IRT) has recently been applied to LLM evaluation (lalor2024item, ; guinet2025evaluating, ). It provides item parameters such as difficulty, discrimination, and guessing, as well as latent ability estimates $\theta$ for models. However, existing IRT applications remain largely static: tinyBenchmarks (polo2024tinybenchmarks, ) uses clustering for item selection but doesn’t guarantee informativeness for $\theta$ estimation, while MetaBench (metabench, ) requires computationally expensive iterations to identify stable subsets. These approaches often lack proper psychometric validation and emphasize predictive accuracy over model fit diagnostics, which makes it difficult to ensure that ability estimates are valid, interpretable, and comparable across models.
Beyond these limitations of existing IRT applications, many evaluations continue to rely on average scores. Average scores tend to mask meaningful model differences and are often affected by form-dependence, nonlinear scaling, equal weighting of uninformative items, and contamination sensitivity (see Appendix A for detailed analysis). In contrast, IRT-based ability estimates ($\theta$) provide form-invariant, uncertainty-aware alternatives that adjust for item difficulty and discrimination.
2.2 Adaptive Testing
Computerized adaptive testing (CAT) is a psychometric framework for efficiently measuring ability by selecting items that maximize Fisher information. By adapting item selection in real time, CAT caters to different ability levels, administering easier items to weaker models and more challenging items to stronger ones. CAT can operate with a fixed test length or with a variable length, terminating once a predefined precision threshold is reached (weiss1982improving, ).
The challenge of evaluating LLMs closely parallels that of international large-scale assessments for humans. Our dataset includes more than 4,000 models, ranging from simple pattern-matching systems with very low ability estimates ($\theta<-2$) to advanced models such as GPT-4 with high ability estimates ($\theta>2$) . This wide heterogeneity makes static benchmarks inadequate: fixed test forms inevitably include items that are trivial for advanced models and impossible for weaker ones, which limits their diagnostic value. Unlike static benchmarks, CAT explicitly adapts to each model’s ability level, ensuring that item selection remains informative across the full spectrum of performance.
CAT allows us to reduce test length while maintaining measurement precision. Modern extensions such as Multi-Stage Testing and the incorporation of process data (zenisky2009multistage, ; zheng2015fly, ; tang2024utilizing, ) provide further gains in efficiency and validity. Prior attempts to explore adaptive evaluation for LLMs have been either limited in scope (zhuang2023efficiently, ) or primarily conceptual (zhuangposition, ). Beyond efficiency, CAT also strengthens defenses against contamination. Because models are evaluated on small, adaptively chosen subsets selected in real time, the overlap between training and evaluation data is substantially reduced. Furthermore, CAT protocols can be refreshed with new items while preserving comparability of ability estimates $\theta$. A detailed comparison of IRT-based approaches is provided in Appendix B.
3 Methodology
We introduce a novel adaptive testing framework that transforms LLM evaluation from static benchmarking to dynamic ability estimation. Our approach addresses three critical limitations of current evaluation practice: (1) it distinguishes models with identical average scores but different capability patterns, (2) it reduces computational cost by requiring 90% fewer items while maintaining accuracy, and (3) it naturally mitigates contamination risks through diverse item sampling.
This section presents our framework in four stages: problem formulation (Section 3.1), data construction with psychometric filtering (Section 3.2), item bank calibration using IRT models (Section 3.3), and adaptive testing with randomesque selection (Section 3.4).
3.1 Problem Formulation and Setup
We formulate LLM evaluation as a psychometric measurement problem. Let $\mathcal{I}$ denote the set of benchmark items and $\mathcal{L}$ the set of language models. For each model $\ell\in\mathcal{L}$ and item $i\in\mathcal{I}$, we observe a binary response $Y_{i,\ell}\in\{0,1\}$, where 1 indicates correct and 0 incorrect. These responses form the item-response matrix $\{Y_{i,\ell}\}_{i\in\mathcal{I},\ell\in\mathcal{L}}$.
Unlike traditional approaches that rely solely on accuracy scores, our objective is to estimate the latent ability $\theta_{\ell}$ of each model based on its response pattern $\{Y_{i,\ell}\}_{i\in\mathcal{I}}$, while simultaneously calibrating item-level parameters: discrimination $a_{i}$, difficulty $b_{i}$, and guessing $c_{i}$. This approach enables fine-grained model comparison even when models achieve identical accuracy, as $\theta_{\ell}$ accounts for the varying informativeness of different items.
3.2 Discriminative Item Filtering
We construct the item-response matrix using data from the HuggingFace Open LLM Leaderboard. The item pool $\mathcal{I}$ spans five benchmarks: ARC, GSM8K, HellaSwag, TruthfulQA, and WinoGrande. To ensure data quality for IRT calibration, we apply two levels of filtering: removing unsuitable models and eliminating non-informative items.
Model Selection. We retain only models $\mathcal{L}$ with complete responses across all items to ensure unbiased ability estimates. Models with extreme scores (below 0.1st percentile) are excluded to prevent parameter estimation instability, as IRT’s sigmoidal functions become under-constrained at the boundaries.
Item Filtering. We apply two complementary filters to retain only discriminative items:
  • Low-variance removal: Items with response standard deviation $1\%$ or mean accuracy $95\%$ are discarded, as they provide litter information for differentiating between models.
  • Discrimination filtering: We compute the point-biserial correlation $r_{pb}(i)$ between each item’s response vector $\{Y_{i,\ell}\}_{\ell\in\mathcal{L}}$ and the models’ total scores $T_{\ell}=\sum_{j\in\mathcal{I}}Y_{j,\ell}$ (see Appendix E for details). Items with $r_{pb}(i)<0.1$ are removed as non-diagnostic.
This filtering process yields a refined response matrix that supports stable and reliable IRT calibration (see Table 5 in Appendix E for detailed results).
3.3 Scalable IRT Calibration
The calibration stage estimates item parameters $(a_{i},b_{i},c_{i})$ and computes reference ability estimates $\hat{\theta}_{\ell}^{\mathrm{whole}}$ for each LLM $\ell$ for validation. To model the probability of a correct response, we adopt the three-parameter logistic (3PL) IRT model (birnbaum1968some, ; lord1980, ):
$p_{i}(\theta_{\ell})=c_{i}+\frac{1-c_{i}}{1+\exp(-a_{i}(\theta_{\ell}-b_{i}))}.$ (1)
Here, $a_{i}$ is the discrimination parameter, which determines how sharply item $i$ differentiates between stronger and weaker models. $b_{i}$ is the difficulty parameter, specifying the ability level at which a model has a 50% chance (beyond guessing) of answering item $i$ correctly. $c_{i}$ is the guessing parameter, setting the lower bound on the probability of success due to random guessing. These parameters enable Fisher information-based prioritization of items in our adaptive framework, distinguishing high-quality items from those with low discriminative power.
Common-Person Calibration at Scale. We develop a partition-based calibration strategy that takes advantage of the unique characteristics of LLM benchmarking. Unlike traditional testing where only subsets of items are administered to each examinee, all LLM models respond to all items, enabling a novel application of common-person linking. We partition items into $K$ non-overlapping subsets $\mathcal{I}_{k}$ (each with $|\mathcal{I}_{k}|\geq 100$ items), calibrate each subset independently, then apply mean-sigma transformations (kolen2014, ) using the model population $\mathcal{L}$ as linking anchors. This approach reduces computational complexity from $O(|\mathcal{I}|^{3})$ to $O(K\cdot\max_{k}|\mathcal{I}_{k}|^{3})$ while maintaining calibration accuracy through the redundancy of having all models as linkers (chalmers2012mirt, ).
Heterogeneity-Aware Ability Estimation. The extreme heterogeneity of LLM populations, spanning from random baselines ($\theta\approx-3$) to near-perfect models ($\theta\approx 3$), breaks standard estimation methods. Standard IRT practice relies on Expected A Posteriori (EAP) estimation (bock1982adaptive, ) for ability estimation. However, EAP produces undefined values for all-correct or all-incorrect response patterns, affecting 12% of our model population spanning from random baselines to near-perfect systems. To address this LLM-specific challenge, we employ Weighted Likelihood Estimation (WLE) (warm1989weighted, ) with bias correction term $\frac{J(\theta)}{2I(\theta)}$ where $J(\theta)=\sum_{i}\frac{\partial I_{i}(\theta)}{\partial\theta}$ (warm1989weighted, ). This choice enables finite estimates across the full ability spectrum while maintaining consistency properties (baker2004item, ), crucial for establishing reliable validation baselines in the extreme heterogeneity of LLM evaluation.
Multi-Subset Model Fit Validation. Unlike prior IRT applications to LLM evaluation (polo2024tinybenchmarks, ; metabench, ) that omit model validation, we conduct rigorous psychometric diagnostics to ensure calibration quality. We report the $M_{2}$ statistic (maydeu2015m2, ) with RMSEA indices, establishing that our 3PL models achieve good fit (RMSEA <0.05<0.05) across all benchmarks, as summarized in Appendix D. Our partition-based calibration strategy, which divides the full item bank into non-overlapping subsets (each containing $\geq 100$ items for statistical stability), enables both computational feasibility and robust validation. Since the same set of models $\mathcal{L}$ acts as common persons across all partitions, diagnostic statistics reflect global calibration quality rather than partition-specific artifacts. This multi-subset linking design ensures that model fit metrics capture systematic patterns across the entire item bank, not just localized subsets. This validation is crucial for reliable adaptive testing, as misfitting items would compromise Fisher information calculations and degrade selection accuracy.
3.4 Adaptive Testing with Information Selection
Our proposed ATLAS dynamically selects the most informative items for each model, dramatically reducing the number of items needed while maintaining accuracy. Algorithm 1 presents the complete procedure. The algorithm includes several key design choices tailored to LLM evaluation:
Algorithm 1 Adaptive Testing for Model $\ell$
1:Initialize: $\hat{\theta}_{0}\leftarrow 0$, test record $R_{\ell}\leftarrow\emptyset$, $t\leftarrow 0$
2:while $t<\mathrm{max\_items}$ and not converged do
3:  $t\leftarrow t+1$
4:  if $t=1$ then
5:   Select item $i_{t}$ with $|b_{i_{t}}-\hat{\theta}_{0}|$ minimized
6:  else
7:   Compute Fisher information $I_{i}(\hat{\theta}_{t-1})$ for all unadministered items
8:   Select $i_{t}$ randomly from top-5 most informative items
9:  end if
10:  Administer item $i_{t}$ to model $\ell$, observe response $Y_{i_{t},\ell}$
11:  Update record: $R_{\ell}\leftarrow R_{\ell}\cup\{(i_{t},Y_{i_{t},\ell})\}$
12:  Update ability: $\hat{\theta}_{t}\leftarrow\mathrm{EAP}(R_{\ell})$
13:  Compute standard error: $\mathrm{SE}(\hat{\theta}_{t})\leftarrow 1/\sqrt{\sum_{j\in R_{\ell}}I_{j}(\hat{\theta}_{t})}$
14:  if $t\geq\mathrm{min\_items}$ and $\mathrm{SE}(\hat{\theta}_{t})\leq\tau$ then
15:   break $\triangleright$ Convergence achieved
16:  end if
17:end while
[…]
Initialization and Bounds. We initialize the ability estimate at $\hat{\theta}_{0}=0$. In IRT, abilities are typically standardized to mean zero, so this choice provides a neutral starting point that avoids bias toward high- or low-performing models and stabilizes early item selection when little information is available. We enforce minimum (30) and maximum (500) item limits. The minimum ensures stable estimation for models at performance extremes, while the maximum constrains computational cost and yields approximately 90% reduction in test length relative to full benchmarks.
Randomesque Item Selection. Rather than deterministically selecting the single most informative item, we randomly sample from the top-5 candidates ranked by Fisher information:
$I_{i}(\theta)=a_{i}^{2}\cdot p_{i}(\theta)\cdot[1-p_{i}(\theta)].$ (2)
This randomesque strategy (kingsbury1989procedures, ) prevents over-reliance on specific item types while still keeping high information, which is important for models with specialized capabilities.
Sequential Ability Updates. After each item administration, we update the ability estimate using Expected A Posteriori (EAP) estimation (bock1982adaptive, ):
$\hat{\theta}_{t}=\mathbb{E}[\theta|R_{\ell}]=\int\theta\cdot p(\theta|R_{\ell})\,d\theta.$ (3)
EAP provides numerically stable updates with sparse early responses and incorporates prior knowledge about ability distributions. In contrast, WLE tends to become unstable when response patterns are extreme, a situation common in the early stages of adaptive testing.
[…]
Output and Validation. For each model $\ell$, the algorithm produces: (1) the administered item sequence and responses $R_{\ell}$, (2) the ability estimate trajectory $\{\hat{\theta}_{t}\}$ with associated standard errors, and (3) the final estimate $\hat{\theta}_{\ell}$. We validate these adaptive estimates against whole-bank references $\hat{\theta}_{\ell}^{\mathrm{whole}}$ to confirm that our dramatic reduction in items does not compromise measurement accuracy.
4 Experiments
We evaluate the proposed ATLAS framework across five benchmarks, comparing its efficiency and accuracy against static baselines while analyzing exposure rates, overlap patterns, and the advantages of IRT-based ability estimation over raw accuracy scores.
4.1 Experimental Setup and Metrics
We evaluate ATLAS across five diverse benchmarks covering different cognitive domains: WinoGrande (commonsense reasoning), TruthfulQA (factual consistency), HellaSwag (procedural inference), GSM8K (mathematical reasoning), and ARC (scientific question answering). All experiments use calibrated item banks from Section 3.3.
We compare against four static baselines that do not adapt to individual models: (1) Random sampling of 100 items uniformly from the full bank, (2) TinyBenchmarks (polo2024tinybenchmarks, ) using predetermined subsets selected via clustering without Fisher information optimization, (3) MetaBench-Primary and (4) MetaBench-Secondary (metabench, ) using curated splits that require computationally expensive iterations to identify stable subsets. Unlike these static approaches, ATLAS uses three precision thresholds ($\mathrm{SE}(\hat{\theta})\leq 0.1,0.2,0.3$) with bounds of 30–500 items, terminating when sufficient precision is achieved.
We evaluate using five metrics: (1) Mean Absolute Error (MAE) between ATLAS estimates $\hat{\theta}_{\ell}$ and full-bank references $\hat{\theta}_{\ell}^{\text{whole}}$ to assess accuracy, (2) Average Number of Items Administered to measure efficiency, (3) Item Exposure Rate quantifying how frequently each item appears across evaluations, (4) Test Overlap Rate $\bar{Q}$ measuring the average proportion of shared items between any two model evaluations, and (5) Runtime per session to assess computational efficiency. The exposure rate is calculated as the percentage of models that encounter each specific item, while the overlap rate represents the average Jaccard similarity between item sets administered to different models. Additional experimental configurations and psychometric interpretations are provided in Appendix C.
Table 1: Baseline and adaptive evaluation results (MAE and Avg. Item). Lower is better $\downarrow$.
Method WinoGrande TruthfulQA HellaSwag GSM8K ARC
MAE Items MAE Items MAE Items MAE Items MAE Items
Random 0.176 100 0.105 100 0.231 100 0.164 100 0.147 100
TinyBenchmarks 0.247 100 0.161 100 0.299 100 0.173 100 0.165 100
MetaBench-P 0.148 133 0.081 154 0.164 93 0.154 100 0.142 145
MetaBench-S 0.210 106 0.074 136 0.238 58 0.162 100 0.133 100
ATLAS0.1 0.149 78 0.067 51 0.154 39 0.159 73 0.099 77
ATLAS0.2 0.181 39 0.074 30 0.159 30 0.181 39 0.126 36
ATLAS0.3 0.183 32 0.072 30 0.161 30 0.186 32 0.128 30
Table 2: Adaptive evaluation efficiency and diversity. ATLAS maintains low item exposure rates ($10\%$) and moderate test overlap ($16-27\%$), providing natural contamination resistance while achieving fast selection times (<1<1 minute per model). Lower values indicate better performance for all metrics.
Benchmark Method Test Overlap $\downarrow$ Avg. Item $\downarrow$ Avg. Selection $\downarrow$
(Item #) Rate (%) Exposure Rate (%) Time (s)
WinoGrande
ATLASSE≤0.1 18.85 9.28 40.99
ATLASSE≤0.2 17.32 5.07 19.92
ATLASSE≤0.3 20.28 4.24 16.74
TruthfulQA
ATLASSE≤0.1 18.93 8.17 15.97
ATLASSE≤0.2 21.69 10.56 9.37
ATLASSE≤0.3 21.67 10.68 9.72
HellaSwag
ATLASSE≤0.1 15.76 3.90 75.52
ATLASSE≤0.2 19.28 6.23 56.93
ATLASSE≤0.3 19.25 6.85 57.06
GSM8K
ATLASSE≤0.1 22.43 8.32 45.69
ATLASSE≤0.2 22.16 5.63 24.08
ATLASSE≤0.3 26.89 8.24 19.06
ARC
ATLASSE≤0.1 19.50 10.08 30.99
ATLASSE≤0.2 19.25 5.42 13.98
ATLASSE≤0.3 21.91 9.77 11.82
4.2 Performance and Reliability Analysis
Figure 1: Comparison of subset (predicted) ability estimates against whole-bank (reference) abilities across five benchmarks. Points along the diagonal indicate perfect agreement. ATLAS maintains the closest alignment overall, particularly on TruthfulQA, ARC, HellaSwag and in the low-ability regime of GSM8K, while static baselines such as TinyBenchmarks and random show greater variance and systematic deviation.
Table 1 reveals ATLAS’s consistent efficiency advantages across all benchmarks. On HellaSwag, ATLAS achieves the best accuracy (0.154 MAE) using only 39 items - just 42% of MetaBench-Primary’s 93 items while improving accuracy by 6%. The efficiency gains are most dramatic on TruthfulQA, where ATLAS achieves the lowest MAE of 0.067 using only 33% of MetaBench-Primary’s items (51 vs 154) and outperforms both MetaBench variants. On ARC, ATLAS demonstrates substantial improvements in both dimensions: 30% better accuracy (0.099 vs 0.142 MAE) using just 53% of the items (77 vs 145). Even on GSM8K where MetaBench-Primary achieves marginally lower error, ATLAS maintains competitive performance (0.159 vs 0.154 MAE) while requiring 27% fewer items. WinoGrande shows similar patterns, with ATLAS matching MetaBench-Primary’s accuracy (0.149 vs 0.148 MAE) using only 59% of the items.
Figure 1 compares ability estimates from subsets against whole-bank references. ATLAS’s tight alignment to the diagonal demonstrates reliable ability estimation across benchmarks. In contrast, static baselines show systematic deviations: TinyBenchmarks exhibits bias at high-ability levels, while random sampling produces broad scatter, confirming the importance of principled item selection over arbitrary subsampling.
4.3 Contamination Resistance and Efficiency Benefits
Table 2 demonstrates ATLAS’s contamination resistance and efficiency advantages across multiple dimensions. The average item exposure rates remain below 10% across all benchmarks and precision levels, with the lowest rates on large banks: HellaSwag achieves just 3.9% exposure with SE $\leq 0.1$, meaning each item appears in fewer than 4% of model evaluations. This contrasts sharply with static baselines where every item is exposed in 100% of evaluations. Even smaller banks maintain low exposure rates, with ARC and TruthfulQA showing 5.4-10.7% exposure rates, making systematic memorization during pretraining practically impossible.
The test overlap rates of 16-27% provide natural contamination resistance while maintaining evaluation consistency. On HellaSwag, the 15.8% overlap rate means two randomly selected models share fewer than 16% of items on average, compared to 100% overlap in static evaluation. This diversity scales with bank size: larger banks like HellaSwag (5,608 items) achieve the lowest overlap rates (15.8-19.3%), while smaller banks like GSM8K show higher but still protective overlap rates (22.2-26.9%). The precision threshold affects this trade-off: stricter thresholds (SE $\leq 0.1$) generally produce lower overlap rates but higher exposure rates due to longer tests.
Runtime efficiency demonstrates practical scalability, with selection times averaging 9.4-75.5 seconds per model across benchmarks. The timing scales roughly with bank size: HellaSwag requires the longest selection time (57-76 seconds) due to its large item pool, while TruthfulQA achieves the fastest selection (9.4-16.0 seconds). Importantly, these times include the full adaptive selection process and remain well under one minute for most configurations, making ATLAS practical for both interactive evaluation and large-scale benchmarking scenarios.
(a) GSM8K
(b) HellaSwag
Figure 2: Comparison of IRT ability estimates $\hat{\theta}_{\ell}^{\text{whole}}$ with raw accuracy. Left: Ability vs accuracy reveals strong correlation but critical differences at performance extremes where accuracy collapses. Right: Rank comparison shows systematic reordering, with 23% (GSM8K) and 31% (HellaSwag) of models shifting >> 10 positions. IRT separates models with identical accuracies by accounting for which items they solve correctly.
4.4 Distinguishing Low- and High-Performing Models
Figure 2 demonstrates IRT’s key advantage: separating models with similar accuracy scores through ability estimates. Despite strong correlations (0.99 for GSM8K, 0.96 for HellaSwag), systematic differences emerge at performance extremes. In low-performing regimes, accuracy collapses into narrow bands (0.10-0.15) while IRT spans $\theta\approx-3$ to 1-1, distinguishing models that solve easy versus challenging items. In high-performing regimes, ceiling effects compress accuracy differences, but IRT maintains discrimination across $\theta\approx 1.5$ to $2.5$. Similar patterns are observed across TruthfulQA, WinoGrande, and ARC benchmarks, with additional score-versus-theta comparisons provided in Appendix G.
The right panels show systematic rank reordering: 23-31% of models shift >> 10 positions when ranked by IRT versus accuracy. Models performing well on hard items receive higher IRT ranks despite moderate accuracy, while those succeeding on easy items are appropriately downweighted. This enhanced discrimination provides more reliable model comparisons, especially critical in saturated performance regions where accuracy-based evaluation fails. Similar patterns of IRT superiority in distinguishing models are observed across TruthfulQA, WinoGrande, and ARC benchmarks, with additional analysis provided in Appendix G.
4.5 Items Are Not Equally Informative
Figure 3: Two models with identical accuracy (0.833) on WinoGrande receive different ability estimates ($\hat{\theta}_{A}=1.2$ vs $\hat{\theta}_{B}=0.6$). Model A succeeds on harder items (darker cells on right), while Model B answers easier items (darker cells on left). IRT captures these item difficulty patterns that raw accuracy cannot.
Figure 3 demonstrates IRT’s key advantage: models with identical accuracy (0.833) receive different ability estimates ($\hat{\theta}_{A}=1.2$ vs $\hat{\theta}_{B}=0.6$) based on which items they solve correctly. Under the 3PL IRT model, items differ in discrimination $a_{i}$, difficulty $b_{i}$, and guessing $c_{i}$ parameters. Model A succeeds on high-difficulty items ($b_{i}>0.5$) with strong discrimination ($a_{i}>1.5$), yielding 2.3× more Fisher information than Model B, which mainly answers easier, less discriminative items ($b_{i}<-0.5$, $a_{i}<0.8$).
IRT provides automatic quality control by weighting items according to their empirical contribution to distinguishing model abilities. In our calibrated banks, 3.2-5.7% of items exhibit negative discrimination ($a_{i}<0$), indicating systematic flaws where stronger models perform worse. For example, WinoGrande item #247 achieves 0.89 accuracy but $a_{i}=-0.43$ due to exploitable linguistic artifacts. Under raw accuracy, this flawed item contributes equally (weight = 1/N) to all scores, potentially inflating weak models. Under IRT, negative discrimination automatically down-weights its contribution by 82% (effective weight $\propto a_{i}^{2}\approx 0.18$), reducing measurement contamination and providing more reliable ability estimates than accuracy alone. Similar patterns of identical accuracy leading to different ability estimates are observed across ARC, HellaSwag, and other benchmarks, with additional heatmap visualizations provided in Appendix G.
5 Discussion
5.1 Implications for Benchmark Design
Our findings underscore several critical considerations for the design of future large language model (LLM) benchmarks. First, reduced item sets necessitate content balancing to preserve assessment validity. Content balancing entails maintaining proportional representation across distinct skill domains or cognitive competencies to ensure that the benchmark reflects the intended construct rather than overemphasizing specific subskills. Without such balance, benchmarks risk producing biased or unrepresentative evaluations. The proposed adaptive framework can be readily extended to optimize both domain coverage and measurement precision simultaneously (cheng2009maximum, ). Achieving comparable balance using static, reduced item sets would be substantially more difficult and would require extensive manual tuning or domain expertise.
Second, item quality exerts a direct influence on evaluation reliability. Our analysis reveals that 3.2–5.7% of items display negative discrimination parameters, indicating systematic flaws that undermine accuracy-based performance estimates. Although IRT automatically attenuates the influence of such items through probabilistic weighting, many existing benchmarks lack analogous quality control mechanisms. Consequently, poorly constructed or ambiguous items can distort aggregate scores and obscure genuine model differences. Future benchmark design should therefore adopt psychometric validation as standard practice, incorporating procedures such as discrimination screening and content alignment checks to enhance the interpretability and fairness of evaluation outcomes.
Third, item exposure rate and test overlap rate are essential metrics for benchmark robustness and reusability. A low exposure rate ensures that individual items are rarely reused across assessments, reducing the risk of model memorization or contamination of test content. Similarly, a low test overlap rate—the proportion of shared items between distinct benchmark forms—preserves evaluation independence and minimizes redundancy. Adaptive benchmarking intrinsically achieves lower exposure and overlap rates by tailoring item selection to each model’s estimated proficiency, generating diverse yet psychometrically comparable test forms. This adaptivity confers a distinct advantage over static benchmarks, enabling more secure, scalable, and sustainable evaluation practices in iterative model development contexts.
5.2 Limitations and Future Work
Despite notable efficiency gains, our framework has several limitations. Initial calibration relies on a representative model population, which may become outdated as architectures evolve. Nonetheless, adaptive testing supports incremental updates: new model responses can be incorporated to refine item parameters and maintain calibration accuracy over time. The current implementation remains limited to multiple-choice formats, while generative or open-ended tasks require alternative scoring and modeling strategies. Moreover, our framework is unidimensional, assuming a single latent proficiency dimension across all items, whereas LLM capabilities are inherently multidimensional. Future work should extend the framework to multidimensional IRT formulations that jointly model reasoning, factuality, and linguistic ability.
To facilitate broader adoption, we have implemented a modular scoring interface that allows users to define custom evaluation functions (e.g., determining whether a model answers a selected item correctly). Future work will explore online or Bayesian calibration techniques for continuous item updating, multidimensional modeling to capture diverse LLM capabilities, and hybrid adaptive–generative evaluation for open-ended tasks. We also plan to extend the system with interfaces for cross-model benchmarking, enhanced item exposure control, and visualization tools that improve interpretability and diagnostic insight.
6 Conclusion
Adaptive testing replaces exhaustive benchmarking with efficient, contamination-resistant measurement. Our ATLAS framework reduces items by 90% while maintaining precision, separates models with identical accuracies using IRT-based ability estimates, and limits contamination through << 10% item exposure rates. With benchmark saturation and contamination risks increasing, adaptive testing provides a rigorous alternative, applying measurement science principles instead of static averaging.
Appendix A Detailed Analysis of Average Score Limitations
Average score (percent correct) remains the most widely reported metric for evaluating LLMs. While it provides a convenient ordinal indicator for fixed forms, it is a shaky measure of underlying ability.
First, average scores are form-dependent: changing the mix or difficulty of items alters percent correct, even if the model’s true ability is unchanged. Second, the metric has a nonlinear scale: improvements at the extremes (e.g., $98\%\!\to 100\%$) do not reflect the same underlying gain as improvements in the middle (e.g., $50\%\!\to 52\%$). Third, it assumes equal informativeness across items, allowing easy or guessable items to influence the mean as much as highly discriminative ones. Fourth, it is subject to coverage bias: the observed score reflects the content blueprint of the test rather than ability across domains. Fifth, average scores offer no measure of uncertainty, making it unclear whether differences are statistically meaningful. Finally, they are highly sensitive to contamination: memorized items from pretraining can artificially inflate percent correct without reflecting genuine reasoning or generalization.
In contrast, IRT-based ability estimates ($\theta$) provide form-invariant, uncertainty-aware measures that adjust for item difficulty and discrimination. Reporting $\theta\pm\text{SE}(\theta)$ offers a psychometrically principled alternative. For communication purposes, reconstructed percent scores may be shown alongside, but $\theta$ should serve as the primary indicator of model capability.
Appendix B Comparison of IRT-Based Benchmark Methods
Table 3: Comparison of IRT-Based Benchmark Methods for LLMs
Factor TinyBenchmarks (Static) MetaBench (Static) ATLAS
IRT Calibration Required Required Required
Adaptivity None (same items) None (same items) High (items vary by ability)
Test Length Fixed Fixed Variable, stopping rules
Exposure control High (same items reused) High (same items reused) Low (rotating pool)
Pool Sensitivity Subset dependent Subset dependent Robust to large pools
Fairness Biased if mis-targeted Biased if mis-targeted Balanced across abilities
Score Precision Low at extremes Low at extremes High, SEs available
Model Fit Rarely checked Rarely checked Possible fit checks
Efficiency Fixed (100 items) Fixed (100 items) Fewer items, comparable accuracy
Saturation Risk High High Low
Appendix C Detailed Experimental Setup and Metrics
C.1 Benchmarks and Datasets
We conduct experiments on five diverse benchmarks covering different cognitive domains:
  • WinoGrande: Commonsense reasoning with pronoun resolution
  • TruthfulQA: Factual consistency and truthfulness evaluation
  • HellaSwag: Procedural inference and common sense completion
  • GSM8K: Mathematical word problems requiring multi-step reasoning
  • ARC: Scientific question answering across multiple domains
All experiments use the calibrated item banks from Section 3.3, ensuring consistent filtering and parameter quality across datasets. Items with poor psychometric properties (negative discrimination $a_{i}<0$, extreme difficulty $|b_{i}|>4$, or unrealistic guessing $c_{i}>0.5$) were excluded during calibration.
C.2 Baseline Configurations
We compare ATLAS against four fixed, non-adaptive strategies:
Random Baseline: Samples 100 items uniformly from the full bank without consideration of item parameters or model ability.
TinyBenchmarks: Uses the predetermined subset from [8], selected via clustering methods but without explicit Fisher information optimization for ability estimation.
MetaBench-Primary and MetaBench-Secondary: Curated splits from [9] that require computationally expensive iterations to identify stable subsets. These splits emphasize predictive accuracy over psychometric validity.
All baseline data is available on Hugging Face: tinyBenchmarks, HCAI/metabench.
Unlike ATLAS, these approaches do not adapt to individual test-takers and serve only as static reference points for accuracy–efficiency tradeoffs.
C.3 ATLAS Configuration Details
For each model $\ell$, we run ATLAS under three precision-based stopping thresholds:
  • $\mathrm{SE}(\hat{\theta})\leq 0.1$: High precision, suitable for fine-grained model comparison
  • $\mathrm{SE}(\hat{\theta})\leq 0.2$: Moderate precision, balancing accuracy and efficiency
  • $\mathrm{SE}(\hat{\theta})\leq 0.3$: Lower precision, maximizing efficiency for rapid screening
A minimum of 30 items is enforced to prevent premature termination due to lucky guesses or initial high-information items, while the maximum is capped at 500 items to ensure computational feasibility. This setup balances precision and budget constraints, simulating realistic conditions for adaptive evaluation in production environments.
C.4 Detailed Metric Definitions
Average Mean Absolute Error (MAE):
$\text{MAE}=\frac{1}{|\mathcal{M}|}\sum_{\ell\in\mathcal{M}}|\hat{\theta}_{\ell}-\hat{\theta}_{\ell}^{\text{whole}}|$
where $\mathcal{M}$ is the set of evaluated models, $\hat{\theta}_{\ell}$ is the ATLAS-derived ability estimate, and $\hat{\theta}_{\ell}^{\text{whole}}$ is the reference estimate from the complete item bank.
Average Item Exposure Rate:
$\text{Exposure}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\frac{\text{count}(i)}{|\mathcal{M}|}$
where $\text{count}(i)$ is the number of times item $i$ was administered across all models $\mathcal{M}$, and $|\mathcal{I}|$ is the total number of items in the bank.
Test Overlap Rate:
$\bar{Q}=\frac{1}{\binom{|\mathcal{M}|}{2}}\sum_{\ell_{1}<\ell_{2}}\frac{|\mathcal{T}_{\ell_{1}}\cap\mathcal{T}_{\ell_{2}}|}{\min(|\mathcal{T}_{\ell_{1}}|,|\mathcal{T}_{\ell_{2}}|)}$
where $\mathcal{T}_{\ell}$ represents the set of items administered to model $\ell$.
C.5 Psychometric Interpretation Guidelines
Item Exposure Rate Benchmarks: Values below 20–30% are considered desirable in large-scale adaptive testing, as they indicate that no small subset of items dominates the evaluation, reducing security risks and promoting item bank longevity.
Test Overlap Rate Benchmarks: Values under 10–15% are typically considered good in adaptive testing practice, providing sufficient form diversity to resist contamination while maintaining measurement consistency.
Runtime Considerations: All measurements reflect computational efficiency in fully automated settings. Runtime includes item selection via Fisher information maximization but excludes model inference time, which varies significantly across architectures.
Appendix D Model Fit Evaluation across Datasets
Table 4 summarizes the average Root Mean Square Error of Approximation (RMSEA) from the MIRT M2 model fit across multiple benchmark datasets. Overall, all datasets exhibit low RMSEA values (ranging from 0.0436 to 0.0690), indicating a generally good fit between the latent multidimensional item response model and the observed response data. Among the evaluated datasets, GSM8K achieves the lowest average RMSEA (0.0436), suggesting that the underlying latent trait structure aligns particularly well with the response patterns, likely due to its consistent problem-solving format. In contrast, TruthfulQA shows a slightly higher RMSEA (0.0690), reflecting greater variability or multidimensionality in the response space.
The narrow gaps between the 5% and 95% RMSEA bounds across datasets indicate model stability and relatively low dispersion of fit indices among individual files. This consistency across diverse benchmarks such as Winogrande, HellaSwag, and ARC supports the reliability of the MIRT calibration process. Taken together, these results suggest that the applied MIRT M2 model provides a satisfactory and robust representation of item-level relationships across heterogeneous reasoning and language understanding tasks.
Table 4: Summary of MIRT M2 Average RMSEA across datasets
Dataset Avg. RMSEA Avg. 5% RMSEA Avg. 95% RMSEA Num. Files
Winogrande $0.0565$ $0.0561$ $0.0568$ $10$
TruthfulQA $0.0690$ $0.0686$ $0.0693$ 66
HellaSwag $0.0482$ $0.0478$ $0.0486$ $50$
GSM8K $0.0436$ $0.0433$ $0.0440$ $12$
ARC $0.0588$ $0.0585$ $0.0592$ 88
  • Note: ‘Avg. RMSEA‘ is the mean RMSEA across all files for each dataset. ‘Avg. 5% RMSEA‘ and ‘Avg. 95% RMSEA‘ are the averages of the file-level 5th and 95th percentile RMSEA bounds, respectively.
Appendix E Data Preprocessing Details
E.1 Point-Biserial Correlation Formula
The point-biserial correlation [23] for item $i$ is defined as:
$r_{pb}(i)\;=\;\frac{\bar{T}_{\,\ell\mid Y_{i\ell}=1}-\bar{T}_{\,\ell\mid Y_{i\ell}=0}}{s_{T}}\cdot\sqrt{p_{i}q_{i}},$
where $\bar{T}_{\,\ell\mid Y_{i\ell}=1}$ and $\bar{T}_{\,\ell\mid Y_{i\ell}=0}$ are the mean total scores of models that answered item $i$ correctly and incorrectly, respectively; $s_{T}$ is the standard deviation of total scores $\{T_{\ell}\}$; $p_{i}=\tfrac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}Y_{i\ell}$ is the proportion of models that answered item $i$ correctly; $q_{i}=1-p_{i}$; and $|\mathcal{L}|$ is the number of models.
E.2 Filtering Results
WinoGrande TruthfulQA HellaSwag GSM8K ARC
Num. Selected Models 4680 4635 3467 4195 4162
Num. Selected Item 1046 628 5608 1306 842
Avg. $M_{2}$ RMSEA 0.0565 0.0690 0.0482 0.0436 0.0588
Table 5: Summary of model and item filtering per benchmark. RMSEA values reflect average fit from $M_{2}$ statistics under the 3PL IRT model. Calibration and ability estimation were performed using the mirt package in R, with WLE as the estimator. Items were filtered based on variability, ceiling effects, and point-biserial discrimination.
Appendix F Evaluation Metric Definitions
This section provides the exact mathematical definitions of the evaluation metrics introduced in Section 4, along with brief interpretations.
Average Mean Absolute Error (MAE).
For each model $\ell$, let $\hat{\theta}_{\ell}$ denote the CAT-derived ability estimate and $\hat{\theta}_{\ell}^{\text{whole}}$ the full-bank reference estimate. The average MAE is:
$\text{MAE}=\frac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}\left|\hat{\theta}_{\ell}-\hat{\theta}_{\ell}^{\text{whole}}\right|,$ (4)
where $|\mathcal{L}|$ is the number of models evaluated. Interpretation: Lower MAE indicates higher fidelity, i.e., CAT ability estimates more closely match the whole-bank reference.
Average Item Exposure Rate.
Let $h_{i}$ denote the number of models administered item $i$, with $|\mathcal{I}|$ total items and $|\mathcal{L}|$ total models. The item exposure probability for item $i$ is
$P(A_{i})=\frac{h_{i}}{|\mathcal{L}|}.$ (5)
The average item exposure rate is then
$\bar{P}(A_{i})=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}P(A_{i}).$ (6)
Interpretation: Lower values indicate higher adaptivity and greater item diversity, while higher values suggest uniform or repetitive item usage across models.
Test Overlap Rate.
Following chen2005simplified [24], the expected proportion of common items between two randomly selected test forms is given by
$\bar{Q}=\frac{|\mathcal{L}|\sum_{i=1}^{|\mathcal{I}|}P(A_{i})^{2}}{\bar{L}(|\mathcal{L}|-1)}-\frac{1}{|\mathcal{L}|-1},$ (7)
where $\bar{L}$ is the average test length. Interpretation: Lower values of $\bar{Q}$ imply greater test form diversity, which reduces risks of collusion and item memorization.
Correlation Metrics
For completeness, we provide the definitions of the rank-based correlation coefficients used in Section 4.
Spearman correlation.
$\rho=1-\frac{6\sum_{i=1}^{n}d_{i}^{2}}{n(n^{2}-1)},$
where $d_{i}$ is the rank difference for observation $i$ across the two measures.
Kendall correlation.
$\tau=\frac{(\#\text{concordant pairs})-(\#\text{discordant pairs})}{\tfrac{1}{2}n(n-1)}.$
Appendix G Additional Experimental Results
While both TinyBenchmarks [8] and MetaBench [9] include MMLU [25] as part of their evaluation suites, they treat it as a single unified dataset by aggregating all 57 subject areas. In contrast, we perform evaluation on a per-subject basis. This design choice acknowledges the heterogeneous nature of MMLU, where each subject represents a distinct knowledge domain with varying linguistic characteristics, content distributions, and difficulty levels. Aggregating across all subjects can obscure these domain-specific patterns and limit interpretability in adaptive assessment.
Since each MMLU subject contains approximately 150 items, we randomly selected several representative subjects—Anatomy, Astronomy, and High-School Macroeconomics—to illustrate ATLAS’s behavior under data-sparse conditions. The corresponding results are reported in Table 6. Despite the small number of items per subject, ATLAS consistently demonstrates robust adaptive evaluation performance. As the selection threshold is relaxed ($SE\leq 0.1\rightarrow 0.3$), the mean absolute error (MAE) increases moderately (e.g., from 0.099 to 0.235 in Anatomy), while the number of evaluated items is substantially reduced (approximately 80%). This indicates that ATLAS effectively balances efficiency and accuracy, even in limited-data regimes.
Moreover, reductions in test overlap and exposure rates across the evaluated subjects suggest that the adaptive mechanism achieves broader item coverage and mitigates redundancy. Evaluation time also decreases proportionally with the number of items, confirming the computational efficiency of the adaptive process.
Table 6: Adaptive evaluation results on MMLU subsets (item range: 10–100). Lower is better $\downarrow$.
Benchmark Method MAE $\downarrow$ Avg. Item $\downarrow$ Test Overlap $\downarrow$ Exposure Rate $\downarrow$ Avg. Time (s) $\downarrow$
MMLU-Anatomy
(113 items)
ATLASSE≤0.1 0.099 100.00 0.94 0.88 4.93
ATLASSE≤0.2 0.149 53.29 0.54 0.47 2.70
ATLASSE≤0.3 0.235 20.49 0.32 0.18 0.98
MMLU-Astronomy
(136 items)
ATLASSE≤0.1 0.099 93.22 0.82 0.69 6.30
ATLASSE≤0.2 0.157 48.21 0.48 0.35 3.29
ATLASSE≤0.3 0.235 11.80 0.40 0.11 0.79
MMLU-High School
Macroeconomics (281 items)
ATLASSE≤0.1 0.089 87.94 0.48 0.31 10.49
ATLASSE≤0.2 0.197 27.64 0.22 0.13 3.46
ATLASSE≤0.3 0.240 18.26 0.20 0.09 2.34
Figure 4: Comparison of raw average scores and whole-bank ability estimates on TruthfulQA. (Left) While average scores compress performance at the extremes, whole-bank ability estimates reveal clearer separation among both low- and high-performing models, reflecting sensitivity to item difficulty and discrimination. (Right) Rank comparison shows strong consistency between the two measures (Spearman $\rho=0.97$, Kendall $\tau=0.87$), but ability-based ranking provides finer resolution, especially in distinguishing weaker and stronger models beyond what raw accuracy captures.
Figure 5: Comparison of raw average scores and whole-bank ability estimates on WinoGrande. (Left) Whole-bank estimates show a non-linear relationship with average score and reveal clearer separation on high-performing models, highlighting that ability captures relative item difficulty and provides finer differentiation beyond raw accuracy. (Right) Rank comparison indicates strong but imperfect alignment (Spearman $\rho=0.86$, Kendall $\tau=0.70$), with deviations from the diagonal reflecting cases where ability-based ranking distinguishes models more effectively than accuracy alone.
Figure 6: Comparison of raw average scores and whole-bank ability estimates on ARC. (Left) Whole-bank estimates exhibit a non-linear relationship with average scores, providing clearer separation on high-performing models by accounting for item difficulty and discrimination. (Right) Rank comparison shows strong but not perfect alignment between the two metrics (Spearman $\rho=0.91$, Kendall $\tau=0.77$), with deviations from the diagonal highlighting cases where ability-based ranking offers more informative distinctions than raw accuracy alone.
Figure 7: Two models with the similar average accuracy (0.713) and (0.714) on ARC nevertheless receive very different whole-bank ability estimates. Model A (mera-mix-4x7B) attains a whole-bank ability rank of 270 because its correct responses are concentrated on more difficult items. In contrast, Model B (LLaMAAntino-3-ANITA-8B-Inst-DPO-ITA) is assigned a much lower whole-bank ability rank of 2612, as its successes occur primarily on easier items. This divergence shows how IRT-based ability estimation can distinguish models that appear identical under raw accuracy by accounting for item difficulty.
Figure 8: Two models with the same average accuracy (0.853) on HellaSwag nevertheless receive very different whole-bank ability estimates. Model A (supermario_v1) attains a whole-bank ability rank of 347 because its correct responses are concentrated on more difficult items. In contrast, Model B (contaminated_proof_7b_v1.0_safetensor) is assigned a much lower whole-bank ability rank of 3074, as its successes occur primarily on easier items. This divergence shows how IRT-based ability estimation can distinguish models that appear identical under raw accuracy by accounting for item difficulty.

### v2

@@
 Abstract
-Large language model evaluation requires thousands of benchmark items, making evaluations expensive and slow.
-Existing methods compute average accuracy across fixed item sets, treating all items equally despite varying quality and informativeness.
-We present ATLAS  an adaptive testing framework using Item Response Theory (IRT) to estimate model ability through Fisher information-guided item selection.
-Our analysis of five major benchmarks reveals that 3-6% of items exhibit negative discrimination, indicating annotation errors that corrupt static evaluation.
-ATLAS achieves 90% item reduction while maintaining measurement precision: on HellaSwag (5,608 items), we match full-benchmark estimates using only 42 items with 0.154 MAE.
-Our framework maintains item exposure rates below 10% and test overlap at 16-27%, compared to static benchmarks where every model sees all items (100% exposure).
-Among 4,000+ tested models, IRT ranks differ from accuracy ranks: models with the same accuracy get different IRT scores, and 23-31% of all models shift by more than 10 rank positions.
-Code and calibrated item banks are available at https://github.com/Peiyu-Georgia-Li/ATLAS.git.
+Evaluating large language models (LLMs) typically requires thousands of benchmark items, making the process expensive, slow, and increasingly impractical at scale.
+Existing evaluation protocols rely on average accuracy over fixed item sets, treating all items as equally informative despite substantial variation in difficulty and discrimination.
+We introduce ATLAS, an adaptive testing framework based on Item Response Theory (IRT) that estimates model ability using Fisher information–guided item selection.
+ATLAS reduces the number of required items by up to 90% while maintaining measurement precision.
+For instance, it matches whole-bank ability estimates using only 41 items (0.157 MAE) on HellaSwag (5,600 items).
+We further reconstruct accuracy from ATLAS’s ability estimates and find that reconstructed accuracies closely match raw accuracies across all five benchmarks, indicating that ability $\theta$ preserves the global performance structure.
+At the same time, $\theta$ provides finer discrimination within accuracy-equivalent models: among more than 3,000 evaluated models, 23–31% shift by more than 10 rank positions, and models with identical accuracies receive meaningfully different ability estimates.
+Code and calibrated item banks available at https://github.com/Peiyu-Georgia-Li/ATLAS.git.
 **footnotetext: These authors contributed equally.
 1 Introduction
-Large language model (LLM) evaluation relies on benchmarks with tens of thousands of test items.
-These benchmarks impose high computational costs and evaluation cycles spanning days or weeks.
-Despite benchmark growth—some proposals exceed 100,000 items—evaluation practice remains static: models receive average accuracy scores across fixed item sets.
-This approach ignores statistical information in large datasets and raises questions about efficiency and validity.
-Current evaluation faces three fundamental limitations.
-First, average scores hide meaningful differences between models with distinct error patterns, particularly for lower-performing models where small ability differences are obscured by measurement noise.
-Second, static benchmarks create vulnerability to data contamination as items leak into pretraining corpora, enabling high scores through memorization rather than genuine capability.
-Third, evaluating complete benchmarks is inefficient, treating poorly discriminative items as equally informative as high-quality questions.
-To address these limitations, we propose ATLAS (Adaptive Testing for LLM Ability Scoring), an adaptive evaluation framework based on computerized adaptive testing (CAT; (lord1980, ; wainer2000computerized, ; weiss1982improving, )).
-ATLAS first calibrates benchmark items using three-parameter logistic (3PL) IRT models to estimate item difficulty, discrimination, and guessing parameters (birnbaum1968some, ; hambleton1991fundamentals, ).
+Large language model evaluation relies on benchmarks with tens of thousands of items, which are costly to run and often take days or weeks to complete.
+Even with benchmarks exceeding 100,000 items, evaluation still depends on average accuracy over fixed item sets.
+This practice overlooks valuable statistical information and raises concerns about efficiency and validity.
+Current evaluation practices face three fundamental limitations.
+First, average accuracy over fixed item sets obscure meaningful differences between models with distinct error patterns, especially among lower-performing models where small ability differences are dominated by measurement noise.
+Second, static evaluations treat poorly discriminative items as equally informative as high-quality questions, leading to unreliable and often misleading comparisons.
+Third, evaluating complete benchmarks is inefficient and time-consuming, requiring models to answer hundreds or thousands of items regardless of how much additional information those items provide.
+To address these limitations, we propose ATLAS (Adaptive Testing for LLM Ability Scoring), an adaptive evaluation framework based on computerized adaptive testing (CAT) (Lord, 2012; Wainer et al., 2000; Weiss, 1982).
+ATLAS first calibrates benchmark items using three-parameter logistic (3PL) IRT models to estimate item difficulty, discrimination, and guessing parameters (Birnbaum, 1968; Hambleton et al., 1991).
 Then, rather than administering fixed item sets, ATLAS dynamically selects items with maximum Fisher information for each model’s current estimated ability, terminating when precision thresholds are reached.
-This approach directly addresses all three limitations: Fisher information-guided selection provides precise ability estimates that distinguish models with identical accuracy, dynamic item selection reduces contamination through diverse test forms, and adaptive termination achieves reliable evaluation with significantly fewer items.
-We validate our framework across five major benchmarks (WinoGrande, TruthfulQA, HellaSwag, GSM8K, ARC).
-ATLAS achieves competitive or superior performance to static baselines: on TruthfulQA, it attains the lowest MAE (0.067) using only 51 items; on HellaSwag, it matches MetaBench accuracy (0.154 MAE) while requiring 3× fewer items (39 vs 93).
-Across benchmarks, ATLAS requires 30-78 items on average compared to thousands in full evaluation, while IRT-based estimates reveal systematic rank reordering that exposes limitations of accuracy-based evaluation.
-Our contributions are: (1) We demonstrate fundamental limitations of average-score evaluation and establish the benefits of psychometric ability estimates for LLM evaluation.
-(2) We develop ATLAS, the first large-scale adaptive testing framework for LLMs, achieving 90% item reduction while maintaining measurement precision and natural contamination resistance through $10\%$ item exposure rates.
-(3) We provide comprehensive psychometric analysis of five major benchmarks, revealing that IRT-based estimates cause systematic rank reordering ($23-31\%$ of models shift more than 10 positions) and identifying that $3-6\%$ of items exhibit negative discrimination due to annotation errors, with calibrated item banks released for reproducible evaluation.
+This approach directly addresses all three limitations: Fisher information-guided selection provides precise ability estimates that distinguish models with identical accuracy, dynamic item selection prioritizes highly discriminative items rather than treating all questions as equally informative, and adaptive termination enables reliable evaluation with far fewer items and substantially less time than full-benchmark scoring.
+We evaluate ATLAS across five major benchmarks, including WinoGrande (Sakaguchi et al., 2021), TruthfulQA (Lin et al., 2022), HellaSwag (Zellers et al., 2019), GSM8K (Cobbe et al., 2021), ARC (Clark et al., 2018) and find that it matches or exceeds the accuracy of strong static baselines while using far fewer items.
+For example, ATLAS achieves the lowest MAE on TruthfulQA ($0.064$ with 48 items) and HellaSwag ($0.157$ with 41 items), and matches MetaBench (Kipnis et al., 2025) on WinoGrande while using 2× fewer items (70 vs.
+133).
+It also outperforms TinyBenchmarks (Polo et al., 2024), which uses 97–100 items but yields higher error across all benchmarks.
+Overall, ATLAS requires only 30–89 items per benchmark compared to hundreds in static subsets, and attains the lowest Information Efficiency Score (IES) across all benchmarks, demonstrating the strongest accuracy–efficiency tradeoff.
+Our contributions are: (1) We identify fundamental limitations of average-score evaluation and show that psychometric ability estimates provide more robust and informative comparisons of LLM performance.
+(2) We introduce ATLAS, a large-scale adaptive testing framework for LLMs that achieves up to 90% item reduction while maintaining measurement precision through SE-controlled stopping, enabling flexible and precision-targeted evaluation beyond fixed-length designs.
+(3) We conduct a comprehensive psychometric analysis of five major benchmarks, revealing that IRT-based ability estimation induces substantial rank reordering (23–31% of models shift by more than 10 positions).
+(4) We highlight the importance of rigorous psychometric validation by reporting model-fit statistics (e.g., RMSEA via the M2 statistic) and demonstrating the use of common-person linking to align item parameters efficiently and ensure cross-model comparability.
 2 Related Work
-2.1 Background
-Large-scale benchmarks such as WinoGrande, TruthfulQA, HellaSwag, GSM8K and ARC have become standard for LLM evaluation but impose significant computational costs.
-These static approaches suffer from benchmark saturation, data contamination, and item redundancy, raising concerns about their ability to reflect genuine model capabilities.
-Item Response Theory (IRT) has recently been applied to LLM evaluation (lalor2024item, ; guinet2025evaluating, ).
+2.1 IRT-based Approaches
+Item Response Theory (IRT) has recently been applied to LLM evaluation (Lalor et al., 2024).
 It provides item parameters such as difficulty, discrimination, and guessing, as well as latent ability estimates $\theta$ for models.
-However, existing IRT applications remain largely static: tinyBenchmarks (polo2024tinybenchmarks, ) uses clustering for item selection but doesn’t guarantee informativeness for $\theta$ estimation, while MetaBench (metabench, ) requires computationally expensive iterations to identify stable subsets.
-These approaches often lack proper psychometric validation and emphasize predictive accuracy over model fit diagnostics, which makes it difficult to ensure that ability estimates are valid, interpretable, and comparable across models.
+However, existing IRT applications remain largely static in nature.
+For instance, TinyBenchmarks (Polo et al., 2024) uses clustering for item selection but doesn’t guarantee informativeness for $\theta$ estimation, while MetaBench (Kipnis et al., 2025) requires computationally expensive iterations to identify stable subsets.
+Moreover, these approaches often lack proper psychometric validation and emphasize predictive accuracy over model fit.
+TinyBenchmarks and MetaBench do not report fit statistics.
+Instead, we compute these metrics using their released IRT code (as shown in Table 1).
+This limitation makes it difficult to ensure that the resulting ability estimates are valid, interpretable, and comparable across models.
+A detailed comparison of IRT-based approaches is provided in Appendix B.
 Beyond these limitations of existing IRT applications, many evaluations continue to rely on average scores.
@@
 2.2 Adaptive Testing
-Computerized adaptive testing (CAT) is a psychometric framework for efficiently measuring ability by selecting items that maximize Fisher information.
-By adapting item selection in real time, CAT caters to different ability levels, administering easier items to weaker models and more challenging items to stronger ones.
-CAT can operate with a fixed test length or with a variable length, terminating once a predefined precision threshold is reached (weiss1982improving, ).
-The challenge of evaluating LLMs closely parallels that of international large-scale assessments for humans.
-Our dataset includes more than 4,000 models, ranging from simple pattern-matching systems with very low ability estimates ($\theta<-2$) to advanced models such as GPT-4 with high ability estimates ($\theta>2$) .
-This wide heterogeneity makes static benchmarks inadequate: fixed test forms inevitably include items that are trivial for advanced models and impossible for weaker ones, which limits their diagnostic value.
-Unlike static benchmarks, CAT explicitly adapts to each model’s ability level, ensuring that item selection remains informative across the full spectrum of performance.
-CAT allows us to reduce test length while maintaining measurement precision.
-Modern extensions such as Multi-Stage Testing and the incorporation of process data (zenisky2009multistage, ; zheng2015fly, ; tang2024utilizing, ) provide further gains in efficiency and validity.
-Prior attempts to explore adaptive evaluation for LLMs have been either limited in scope (zhuang2023efficiently, ) or primarily conceptual (zhuangposition, ).
-Beyond efficiency, CAT also strengthens defenses against contamination.
-Because models are evaluated on small, adaptively chosen subsets selected in real time, the overlap between training and evaluation data is substantially reduced.
-Furthermore, CAT protocols can be refreshed with new items while preserving comparability of ability estimates $\theta$.
-A detailed comparison of IRT-based approaches is provided in Appendix B.
+Computerized adaptive testing (CAT) adjusts item administration based on an examinee’s evolving ability estimate (Meijer and Nering, 1999; Van der Linden and Glas, 2010).
+After each response, the test updates the ability estimate and selects the next item using an algorithm that aims to provide the most informative measurement while satisfying test constraints (Weiss, 1982; Chang, 2015; Cheng and Chang, 2009).
+Related adaptive frameworks such as multistage testing and process-data–based approaches apply similar principles and offer additional flexibility and diagnostic information (Zenisky et al., 2009; Zheng and Chang, 2015; Tang et al., 2025).
+These features allow CAT to evaluate examinees efficiently while maintaining rigorous measurement precision.
+This adaptive structure aligns well with challenges in evaluating LLMs, which vary widely in their performance levels.
+Current evaluations often use large static benchmarks in which every model must answer all items, even when many items provide little information about its ability.
+These benchmarks also rarely report empirical item characteristics, so their difficulty range and informativeness across models remain unclear.
+CAT addresses these limitations by selecting items targeted to each model’s estimated ability, which yields more precise and efficient evaluation with far fewer items.
+However, only a few studies have explored adaptive evaluation for LLMs.
+Early efforts were either limited in scope (Zhuang et al., 2023) or primarily conceptual (Zhuang et al., 2025).
+A recent study that is closely related to our work is Fluid Benchmarking (Hofmann et al., 2025), which appeared around the same time as this research.
+Fluid also applies CAT principles to increase evaluation efficiency and models LLM performance on a latent ability scale.
+The two approaches are complementary and differ in several ways.
+Fluid focuses on adaptive evaluation during LLM pretraining, whereas our work examines post-training evaluation.
+Fluid calibrates its IRT model on 102 LMs, while ATLAS uses a substantially larger and more diverse pool of 3,000+ LMs.
+Fluid adopts a fixed-length adaptive design, while ATLAS uses a precision-based stopping rule that terminates the test once the uncertainty of the ability estimate falls below a predefined threshold.
+Precision-based stopping ensures consistent measurement precision across models and avoids administering unnecessary items.
+In addition, we report model-fit statistics to ensure the adequacy of the IRT model before running CAT and provide a transparent description of calibration and linking procedures used to estimate item parameters.
+Both studies demonstrate that adaptive testing methods can be used at different stages of LLM development and under different design choices, which illustrates the broader potential of CAT-based approaches for scalable and precise LLM assessment.
 3 Methodology
 We introduce a novel adaptive testing framework that transforms LLM evaluation from static benchmarking to dynamic ability estimation.
-Our approach addresses three critical limitations of current evaluation practice: (1) it distinguishes models with identical average scores but different capability patterns, (2) it reduces computational cost by requiring 90% fewer items while maintaining accuracy, and (3) it naturally mitigates contamination risks through diverse item sampling.
+Our approach addresses three critical limitations of current evaluation practice: (1) it reduces computational cost by requiring 90% fewer items while maintaining overall performance, (2) it overcomes the ceiling effects of accuracy-based metrics and preserves discrimination across the ability spectrum, and (3) it distinguishes models with identical accuracy score but different abilities.
 This section presents our framework in four stages: problem formulation (Section 3.1), data construction with psychometric filtering (Section 3.2), item bank calibration using IRT models (Section 3.3), and adaptive testing with randomesque selection (Section 3.4).
@@
 This approach enables fine-grained model comparison even when models achieve identical accuracy, as $\theta_{\ell}$ accounts for the varying informativeness of different items.
-3.2 Discriminative Item Filtering
+3.2 Data Construction with psychometric filtering
 We construct the item-response matrix using data from the HuggingFace Open LLM Leaderboard.
@@
 To ensure data quality for IRT calibration, we apply two levels of filtering: removing unsuitable models and eliminating non-informative items.
-Model Selection.
-We retain only models $\mathcal{L}$ with complete responses across all items to ensure unbiased ability estimates.
-Models with extreme scores (below 0.1st percentile) are excluded to prevent parameter estimation instability, as IRT’s sigmoidal functions become under-constrained at the boundaries.
-Item Filtering. We apply two complementary filters to retain only discriminative items:
-  • Low-variance removal: Items with response standard deviation $1\%$ or mean accuracy $95\%$ are discarded, as they provide litter information for differentiating between models.
-  • Discrimination filtering: We compute the point-biserial correlation $r_{pb}(i)$ between each item’s response vector $\{Y_{i,\ell}\}_{\ell\in\mathcal{L}}$ and the models’ total scores $T_{\ell}=\sum_{j\in\mathcal{I}}Y_{j,\ell}$ (see Appendix E for details).
+Model Selection and Splitting.
+We retain only models $\mathcal{L}$ with complete responses across all items.
+To obtain a calibration sample whose ability distribution approximates a Gaussian, a standard assumption for stable IRT estimation, we exclude models in the extreme low-ability tail (below the 0.1st percentile), whose near-zero response patterns destabilize 3PL parameter estimation.
+The high-ability tail is small and non-degenerate, so these models are retained.
+The selected models are then split into training and testing sets using stratified random sampling (10 bins) to ensure that both splits share a similar ability distribution.
+We allocate 90% of the models to the training set for item calibration, and use the remaining 10% as the testing set for evaluating performance in our experiments (see Table 5).
+Item Filtering.
+We apply two complementary filters to retain only discriminative items:
+  • Low-variance removal: Items with response standard deviation $1\%$ or mean accuracy $95\%$ are discarded, as they provide little information for differentiating between models.
+  • Discrimination filtering: We compute the point-biserial correlation $r_{pb}(i)$ between each item’s response vector $\{Y_{i,\ell}\}_{\ell\in\mathcal{L}}$ and the models’ total scores $T_{\ell}=\sum_{j\in\mathcal{I}}Y_{j,\ell}$ (see Appendix C for details).
 Items with $r_{pb}(i)<0.1$ are removed as non-diagnostic.
-This filtering process yields a refined response matrix that supports stable and reliable IRT calibration (see Table 5 in Appendix E for detailed results).
+This filtering process yields a refined response matrix that supports stable and reliable IRT calibration (see Table 5 for detailed results).
 3.3 Scalable IRT Calibration
 The calibration stage estimates item parameters $(a_{i},b_{i},c_{i})$ and computes reference ability estimates $\hat{\theta}_{\ell}^{\mathrm{whole}}$ for each LLM $\ell$ for validation.
-To model the probability of a correct response, we adopt the three-parameter logistic (3PL) IRT model (birnbaum1968some, ; lord1980, ):
+To model the probability of a correct response, we adopt the three-parameter logistic (3PL) IRT model (Birnbaum, 1968; Lord, 2012):
 $p_{i}(\theta_{\ell})=c_{i}+\frac{1-c_{i}}{1+\exp(-a_{i}(\theta_{\ell}-b_{i}))}.$ (1)
@@
 Common-Person Calibration at Scale.
-We develop a partition-based calibration strategy that takes advantage of the unique characteristics of LLM benchmarking.
-Unlike traditional testing where only subsets of items are administered to each examinee, all LLM models respond to all items, enabling a novel application of common-person linking.
-We partition items into $K$ non-overlapping subsets $\mathcal{I}_{k}$ (each with $|\mathcal{I}_{k}|\geq 100$ items), calibrate each subset independently, then apply mean-sigma transformations (kolen2014, ) using the model population $\mathcal{L}$ as linking anchors.
-This approach reduces computational complexity from $O(|\mathcal{I}|^{3})$ to $O(K\cdot\max_{k}|\mathcal{I}_{k}|^{3})$ while maintaining calibration accuracy through the redundancy of having all models as linkers (chalmers2012mirt, ).
+To estimate item characteristics efficiently using the 3PL model, we adopted a partition-based calibration procedure that leverages the unique structure of LLM benchmarking.
+Instead of fitting the full 3PL model to the entire item pool at once, which would be computationally prohibitive, we divided the items into $K$ non-overlapping subsets $\mathcal{I}_{k}$ (each with $|\mathcal{I}_{k}|\geq 100$ items), and calibrated each subset independently.
+This yields multiple provisional difficulty scales that must be aligned.
+Because all models answer all items, the model population serves as a natural set of common persons, allowing us to link the independently calibrated subsets onto a unified scale using mean–sigma transformations (Brennan, 2014).
+This approach reduces computational complexity from $O(|\mathcal{I}|^{3})$ to $O(K\cdot\max_{k}|\mathcal{I}_{k}|^{3})$ while maintaining calibration accuracy due to the stability provided by having all models serve as linking anchors (Chalmers, 2012).
+A detailed description is provided in Appendix C.2.
 Heterogeneity-Aware Ability Estimation.
-The extreme heterogeneity of LLM populations, spanning from random baselines ($\theta\approx-3$) to near-perfect models ($\theta\approx 3$), breaks standard estimation methods.
-Standard IRT practice relies on Expected A Posteriori (EAP) estimation (bock1982adaptive, ) for ability estimation.
-However, EAP produces undefined values for all-correct or all-incorrect response patterns, affecting 12% of our model population spanning from random baselines to near-perfect systems.
-To address this LLM-specific challenge, we employ Weighted Likelihood Estimation (WLE) (warm1989weighted, ) with bias correction term $\frac{J(\theta)}{2I(\theta)}$ where $J(\theta)=\sum_{i}\frac{\partial I_{i}(\theta)}{\partial\theta}$ (warm1989weighted, ).
-This choice enables finite estimates across the full ability spectrum while maintaining consistency properties (baker2004item, ), crucial for establishing reliable validation baselines in the extreme heterogeneity of LLM evaluation.
+LLM populations exhibit extreme heterogeneity, ranging from near-random models ($\theta\approx-3$) to highly capable systems ($\theta\approx 3$).
+To obtain stable and unbiased estimates across this wide ability spectrum, we adopt the Weighted Likelihood Estimator (WLE) (Warm, 1989), which incorporates a bias-correction term $\frac{J(\theta)}{2I(\theta)}$, where $J(\theta)=\sum_{i}\frac{\partial I_{i}(\theta)}{\partial\theta}$.
+WLE provides finite, well-behaved estimates even at ability extremes and maintains desirable consistency properties (Baker and Kim, 2004).
+These characteristics are essential for establishing reliable evaluation baselines under the substantial heterogeneity present in modern LLM benchmarks.
 Multi-Subset Model Fit Validation.
-Unlike prior IRT applications to LLM evaluation (polo2024tinybenchmarks, ; metabench, ) that omit model validation, we conduct rigorous psychometric diagnostics to ensure calibration quality.
-We report the $M_{2}$ statistic (maydeu2015m2, ) with RMSEA indices, establishing that our 3PL models achieve good fit (RMSEA <0.05<0.05) across all benchmarks, as summarized in Appendix D.
-Our partition-based calibration strategy, which divides the full item bank into non-overlapping subsets (each containing $\geq 100$ items for statistical stability), enables both computational feasibility and robust validation.
-Since the same set of models $\mathcal{L}$ acts as common persons across all partitions, diagnostic statistics reflect global calibration quality rather than partition-specific artifacts.
-This multi-subset linking design ensures that model fit metrics capture systematic patterns across the entire item bank, not just localized subsets.
-This validation is crucial for reliable adaptive testing, as misfitting items would compromise Fisher information calculations and degrade selection accuracy.
+Unlike prior IRT-based LLM benchmarks (Polo et al., 2024; Kipnis et al., 2025), which do not report model-fit diagnostics, we perform explicit psychometric validation to ensure calibration quality.
+We compute the limited-information $M_{2}$ statistic with RMSEA (Maydeu-Olivares, 2013) for TinyBenchmarks, MetaBench, and ATLAS (Table 1).
+TinyBenchmarks exhibits severe misfit across all datasets, consistent with its underidentified calibration setting—estimating up to 15 latent traits from only 395 models.
+MetaBench performs better (RMSEA 0.04–0.14) but still shows poor fit on TruthfulQA and marginal fit on ARC, indicating uneven generalization across domains.
+In contrast, ATLAS consistently achieves acceptable or good fit across all benchmarks, demonstrating stable and robust calibration.
+Our partition-based calibration further supports reliable validation by dividing the item bank into non-overlapping subsets (each $\geq$100 items) while using a shared model set as common persons.
+This linking strategy ensures that fit statistics reflect global calibration quality rather than subset-specific artifacts, which is essential for accurate Fisher-information–based adaptive testing.
+Method Winogrande TruthfulQA HellaSwag GSM8K ARC
+RMSEA Fit RMSEA Fit RMSEA Fit RMSEA Fit RMSEA Fit
+TinyBenchmarks 364.24 Poor 371.49 Poor 646.82 Poor 506.60 Poor 369.89 Poor
+MetaBench 0.0524 Accept. 0.1389 Poor 0.0498 Good 0.0423 Good 0.0811 Marginal
+ATLAS 0.0565 Accept. 0.0690 Accept. 0.0482 Good 0.0438 Good 0.0595 Accept.
+Table 1: Model fit comparison across benchmarks using the limited-information statistic $M_{2}$ and its derived Avg.
+RMSEA values.
+Lower Avg.
+RMSEA indicates better model fit.
+Model fit is interpreted according to standard psychometric thresholds: RMSEA << 0.05 = Good fit; 0.05–0.08 = Acceptable fit (abbreviated as Accept.); 0.08–0.10 = Marginal fit; >> 0.10 = Poor fit.
 3.4 Adaptive Testing with Information Selection
@@
 14:  if $t\geq\mathrm{min\_items}$ and $\mathrm{SE}(\hat{\theta}_{t})\leq\tau$ then
-15:   break $\triangleright$ Convergence achieved
+15:   break
 16:  end if
@@
 We initialize the ability estimate at $\hat{\theta}_{0}=0$.
-In IRT, abilities are typically standardized to mean zero, so this choice provides a neutral starting point that avoids bias toward high- or low-performing models and stabilizes early item selection when little information is available.
+This value is a conventional neutral starting point in CAT because the latent trait scale is typically assumed to be centered at zero.
+Beginning at the scale midpoint helps stabilize early item selection by preventing the algorithm from drifting toward artificially high or low values before any response information is available.
 We enforce minimum (30) and maximum (500) item limits.
 The minimum ensures stable estimation for models at performance extremes, while the maximum constrains computational cost and yields approximately 90% reduction in test length relative to full benchmarks.
-Randomesque Item Selection. Rather than deterministically selecting the single most informative item, we randomly sample from the top-5 candidates ranked by Fisher information:
-$I_{i}(\theta)=a_{i}^{2}\cdot p_{i}(\theta)\cdot[1-p_{i}(\theta)].$ (2)
-This randomesque strategy (kingsbury1989procedures, ) prevents over-reliance on specific item types while still keeping high information, which is important for models with specialized capabilities.
-Sequential Ability Updates. After each item administration, we update the ability estimate using Expected A Posteriori (EAP) estimation (bock1982adaptive, ):
+Randomesque Item Selection.
+Rather than deterministically selecting the single most informative item, we randomly sample from the top-5 candidates ranked by Fisher information:
+$I_{i}(\theta)=\frac{\big(p_{i}^{\prime}(\theta)\big)^{2}}{p_{i}(\theta)\,\big(1-p_{i}(\theta)\big)}$ (2)
+This randomesque strategy (Kingsbury and Zara, 1989) prevents over-reliance on specific item types while still keeping high information, which is important for models with specialized capabilities.
+Sequential Ability Updates.
+After each item administration, we update the ability estimate using Expected A Posteriori (EAP) estimation (Bock and Mislevy, 1982):
 $\hat{\theta}_{t}=\mathbb{E}[\theta|R_{\ell}]=\int\theta\cdot p(\theta|R_{\ell})\,d\theta.$ (3)
@@
 In contrast, WLE tends to become unstable when response patterns are extreme, a situation common in the early stages of adaptive testing.
+Precision-Based Stopping.
 […]
@@
 We validate these adaptive estimates against whole-bank references $\hat{\theta}_{\ell}^{\mathrm{whole}}$ to confirm that our dramatic reduction in items does not compromise measurement accuracy.
+(a) ARC
+(b) TruthfulQA
+(c) HellaSwag
+(d) GSM8K
+(e) WinoGrande
+Figure 1: Comparison of subset (predicted) ability estimates against whole-bank (reference) abilities across five benchmarks (graphical illustration complementing Table 2).
+Points along the identity line indicate perfect agreement.
+ATLAS maintains the closest alignment overall, particularly on TruthfulQA, ARC and HellaSwag and in the high-ability regime of WinoGrande, while static baselines such as TinyBenchmarks and Metabench show greater variance and systematic deviation.
+Method WinoGrande TruthfulQA HellaSwag
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Method WinoGrande TruthfulQA HellaSwag
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 4 Experiments
-We evaluate the proposed ATLAS framework across five benchmarks, comparing its efficiency and accuracy against static baselines while analyzing exposure rates, overlap patterns, and the advantages of IRT-based ability estimation over raw accuracy scores.
+We evaluate the proposed ATLAS framework across five benchmarks, comparing its efficiency and accuracy against static baselines such as random sampling, TinyBenchmarks (Polo et al., 2024), MetaBench-Primary and MetaBench-Secondary (Kipnis et al., 2025).
+We report performance-based and efficiency-based metrics, with full metric definitions provided in Appendix E.
 4.1 Experimental Setup and Metrics
@@
 All experiments use calibrated item banks from Section 3.3.
-We compare against four static baselines that do not adapt to individual models: (1) Random sampling of 100 items uniformly from the full bank, (2) TinyBenchmarks (polo2024tinybenchmarks, ) using predetermined subsets selected via clustering without Fisher information optimization, (3) MetaBench-Primary and (4) MetaBench-Secondary (metabench, ) using curated splits that require computationally expensive iterations to identify stable subsets.
-Unlike these static approaches, ATLAS uses three precision thresholds ($\mathrm{SE}(\hat{\theta})\leq 0.1,0.2,0.3$) with bounds of 30–500 items, terminating when sufficient precision is achieved.
-We evaluate using five metrics: (1) Mean Absolute Error (MAE) between ATLAS estimates $\hat{\theta}_{\ell}$ and full-bank references $\hat{\theta}_{\ell}^{\text{whole}}$ to assess accuracy, (2) Average Number of Items Administered to measure efficiency, (3) Item Exposure Rate quantifying how frequently each item appears across evaluations, (4) Test Overlap Rate $\bar{Q}$ measuring the average proportion of shared items between any two model evaluations, and (5) Runtime per session to assess computational efficiency.
-The exposure rate is calculated as the percentage of models that encounter each specific item, while the overlap rate represents the average Jaccard similarity between item sets administered to different models.
-Additional experimental configurations and psychometric interpretations are provided in Appendix C.
-Table 1: Baseline and adaptive evaluation results (MAE and Avg. Item). Lower is better $\downarrow$.
-Method WinoGrande TruthfulQA HellaSwag GSM8K ARC
-MAE Items MAE Items MAE Items MAE Items MAE Items
-Random 0.176 100 0.105 100 0.231 100 0.164 100 0.147 100
-TinyBenchmarks 0.247 100 0.161 100 0.299 100 0.173 100 0.165 100
-MetaBench-P 0.148 133 0.081 154 0.164 93 0.154 100 0.142 145
-MetaBench-S 0.210 106 0.074 136 0.238 58 0.162 100 0.133 100
-ATLAS0.1 0.149 78 0.067 51 0.154 39 0.159 73 0.099 77
-ATLAS0.2 0.181 39 0.074 30 0.159 30 0.181 39 0.126 36
-ATLAS0.3 0.183 32 0.072 30 0.161 30 0.186 32 0.128 30
-Table 2: Adaptive evaluation efficiency and diversity.
-ATLAS maintains low item exposure rates ($10\%$) and moderate test overlap ($16-27\%$), providing natural contamination resistance while achieving fast selection times (<1<1 minute per model).
-Lower values indicate better performance for all metrics.
-Benchmark Method Test Overlap $\downarrow$ Avg. Item $\downarrow$ Avg. Selection $\downarrow$
-(Item #) Rate (%) Exposure Rate (%) Time (s)
-WinoGrande
-ATLASSE≤0.1 18.85 9.28 40.99
-ATLASSE≤0.2 17.32 5.07 19.92
-ATLASSE≤0.3 20.28 4.24 16.74
-TruthfulQA
-ATLASSE≤0.1 18.93 8.17 15.97
-ATLASSE≤0.2 21.69 10.56 9.37
-ATLASSE≤0.3 21.67 10.68 9.72
-HellaSwag
-ATLASSE≤0.1 15.76 3.90 75.52
-ATLASSE≤0.2 19.28 6.23 56.93
-ATLASSE≤0.3 19.25 6.85 57.06
-GSM8K
-ATLASSE≤0.1 22.43 8.32 45.69
-ATLASSE≤0.2 22.16 5.63 24.08
-ATLASSE≤0.3 26.89 8.24 19.06
-ARC
-ATLASSE≤0.1 19.50 10.08 30.99
-ATLASSE≤0.2 19.25 5.42 13.98
-ATLASSE≤0.3 21.91 9.77 11.82
-4.2 Performance and Reliability Analysis
-Figure 1: Comparison of subset (predicted) ability estimates against whole-bank (reference) abilities across five benchmarks.
-Points along the diagonal indicate perfect agreement.
-ATLAS maintains the closest alignment overall, particularly on TruthfulQA, ARC, HellaSwag and in the low-ability regime of GSM8K, while static baselines such as TinyBenchmarks and random show greater variance and systematic deviation.
-Table 1 reveals ATLAS’s consistent efficiency advantages across all benchmarks.
-On HellaSwag, ATLAS achieves the best accuracy (0.154 MAE) using only 39 items - just 42% of MetaBench-Primary’s 93 items while improving accuracy by 6%.
-The efficiency gains are most dramatic on TruthfulQA, where ATLAS achieves the lowest MAE of 0.067 using only 33% of MetaBench-Primary’s items (51 vs 154) and outperforms both MetaBench variants.
-On ARC, ATLAS demonstrates substantial improvements in both dimensions: 30% better accuracy (0.099 vs 0.142 MAE) using just 53% of the items (77 vs 145).
-Even on GSM8K where MetaBench-Primary achieves marginally lower error, ATLAS maintains competitive performance (0.159 vs 0.154 MAE) while requiring 27% fewer items.
-WinoGrande shows similar patterns, with ATLAS matching MetaBench-Primary’s accuracy (0.149 vs 0.148 MAE) using only 59% of the items.
-Figure 1 compares ability estimates from subsets against whole-bank references.
-ATLAS’s tight alignment to the diagonal demonstrates reliable ability estimation across benchmarks.
-In contrast, static baselines show systematic deviations: TinyBenchmarks exhibits bias at high-ability levels, while random sampling produces broad scatter, confirming the importance of principled item selection over arbitrary subsampling.
-4.3 Contamination Resistance and Efficiency Benefits
-Table 2 demonstrates ATLAS’s contamination resistance and efficiency advantages across multiple dimensions.
-The average item exposure rates remain below 10% across all benchmarks and precision levels, with the lowest rates on large banks: HellaSwag achieves just 3.9% exposure with SE $\leq 0.1$, meaning each item appears in fewer than 4% of model evaluations.
-This contrasts sharply with static baselines where every item is exposed in 100% of evaluations.
-Even smaller banks maintain low exposure rates, with ARC and TruthfulQA showing 5.4-10.7% exposure rates, making systematic memorization during pretraining practically impossible.
-The test overlap rates of 16-27% provide natural contamination resistance while maintaining evaluation consistency.
-On HellaSwag, the 15.8% overlap rate means two randomly selected models share fewer than 16% of items on average, compared to 100% overlap in static evaluation.
-This diversity scales with bank size: larger banks like HellaSwag (5,608 items) achieve the lowest overlap rates (15.8-19.3%), while smaller banks like GSM8K show higher but still protective overlap rates (22.2-26.9%).
-The precision threshold affects this trade-off: stricter thresholds (SE $\leq 0.1$) generally produce lower overlap rates but higher exposure rates due to longer tests.
-Runtime efficiency demonstrates practical scalability, with selection times averaging 9.4-75.5 seconds per model across benchmarks.
-The timing scales roughly with bank size: HellaSwag requires the longest selection time (57-76 seconds) due to its large item pool, while TruthfulQA achieves the fastest selection (9.4-16.0 seconds).
-Importantly, these times include the full adaptive selection process and remain well under one minute for most configurations, making ATLAS practical for both interactive evaluation and large-scale benchmarking scenarios.
+We compare against four static baselines that do not adapt to individual models: (1) Random sampling of 100 items uniformly from the full bank, (2) TinyBenchmarks (Polo et al., 2024) using predetermined subsets selected via clustering without Fisher information optimization, (3) MetaBench-Primary and (4) MetaBench-Secondary (Kipnis et al., 2025) using curated splits that require computationally expensive iterations to identify stable subsets.
+Unlike these static approaches, ATLAS uses three precision thresholds ($\mathrm{SE}(\hat{\theta})\leq 0.1,0.2,0.3$) and an item bound of 30–500, terminating adaptively when the required precision is achieved or the maximum test length is reached.
+Additional experimental configurations are provided in Appendix D.
+We evaluate each method primarily in ability space, comparing ATLAS-estimated abilities $\hat{\theta}_{\ell}$ with full-bank abilities $\hat{\theta}^{\text{whole}}_{\ell}$ (See Table 2).
+We report four metrics: (1) Mean Absolute Error (MAE) to measure estimation accuracy; (2) Standard Error (SE) of the absolute errors across models to quantify stability; (3) Average Test Length, the number of items administered per model; and (4) Information Efficiency Score (IES), which jointly reflects accuracy and item usage relative to a 100-item random baseline (values <1<1 indicate higher efficiency).
+More details of these metrics are provided in Appendix E.
+We also provide accuracy-space evaluations, comparing reconstructed accuracies to full-bank raw accuracies (See Table 3).
+Additional evaluation metrics and their results, including Item Exposure Rate, Test Overlap Rate, and Selection Time are provided in Appendix G.2.
 (a) GSM8K
@@
 IRT separates models with identical accuracies by accounting for which items they solve correctly.
-4.4 Distinguishing Low- and High-Performing Models
-Figure 2 demonstrates IRT’s key advantage: separating models with similar accuracy scores through ability estimates.
-Despite strong correlations (0.99 for GSM8K, 0.96 for HellaSwag), systematic differences emerge at performance extremes.
-In low-performing regimes, accuracy collapses into narrow bands (0.10-0.15) while IRT spans $\theta\approx-3$ to 1-1, distinguishing models that solve easy versus challenging items.
-In high-performing regimes, ceiling effects compress accuracy differences, but IRT maintains discrimination across $\theta\approx 1.5$ to $2.5$.
-Similar patterns are observed across TruthfulQA, WinoGrande, and ARC benchmarks, with additional score-versus-theta comparisons provided in Appendix G.
-The right panels show systematic rank reordering: 23-31% of models shift >> 10 positions when ranked by IRT versus accuracy.
-Models performing well on hard items receive higher IRT ranks despite moderate accuracy, while those succeeding on easy items are appropriately downweighted.
-This enhanced discrimination provides more reliable model comparisons, especially critical in saturated performance regions where accuracy-based evaluation fails.
-Similar patterns of IRT superiority in distinguishing models are observed across TruthfulQA, WinoGrande, and ARC benchmarks, with additional analysis provided in Appendix G.
+Figure 3: Two models with identical accuracy (0.833) on WinoGrande receive different ability estimates ($\hat{\theta}_{A}=1.2$ vs $\hat{\theta}_{B}=0.6$).
+Difficulty levels are ordered from easiest (D1) to hardest (D10).
+Model A performs better on harder items (darker cells on right), while Model B primarily succeeds on easier ones (darker cells on left), highlighting how IRT distinguishes latent ability differences that accuracy fails to capture.
+4.2 Performance and Reliability Analysis
+Table 2 presents a comparison of whole-bank ability estimates and subset-based estimates across benchmarks (results for GSM8K and ARC are provided in Appendix G.1).
+ATLAS consistently delivers the strongest accuracy–efficiency tradeoff among all methods.
+Across every benchmark, an ATLAS variant achieves the lowest Information Efficiency Score (IES), indicating that it achieves the best accuracy-efficiency trade-off.
+For example, ATLAS attains the best MAE on TruthfulQA (0.064 with 48 items) and HellaSwag (0.157 with 41 items), and matches the performance of MetaBench-Primary on WinoGrande while requiring nearly half as many items (70 vs.
+133).
+Even in more challenging settings such as GSM8K and ARC (Appendix G.1), ATLAS maintains low MAE with item counts as small as 30–36, outperforming all static baselines in information efficiency.
+In contrast, static subsets such as TinyBenchmarks and MetaBench exhibit inconsistent performance.
+They are strong on some datasets but poor on others, with substantially higher IES values.
+Overall, these results demonstrate that adaptive item selection enables ATLAS to deliver high-fidelity ability estimates while dramatically reducing evaluation cost, achieving reliable performance that fixed subsets fail to match.
+Figure 1 plots subset-based ability estimates against whole-bank references across all five benchmarks.
+Across datasets, ATLAS shows the closest alignment to this identity line, with tightly clustered points and minimal systematic drift, reflecting stable and high-fidelity ability estimation.
+In contrast, static baselines exhibit benchmark-dependent failures.
+Random sampling yields the widest scatter on every benchmark.
+TinyBenchmarks consistently deviate from the identity line in the extreme high or low-ability regime, especially on ARC and TruthfulQA.
+MetaBench-Primary and MetaBench-Secondary performs reasonably on GSM8K but breaks down substantially on HellaSwag, where both its primary and secondary subsets produce large downward deviations, indicating poor item coverage.
+As shown in Table 8, ATLAS produces adaptive tests with low redundancy and high efficiency.
+Test overlap rates remain between 11–23%, and item exposure stays below 12% across all benchmarks, reflecting broad utilization of the item bank.
+End-to-end runtimes range from 9.4 to 75.5 seconds per model and scale consistently with bank size (shortest for TruthfulQA and longest for HellaSwag), confirming that ATLAS is computationally practical.
+Accuracy Reconstruction.
+To assess whether the estimated ability $\theta$ is consistent with traditional accuracy-based evaluation, we reconstruct each model’s expected percent correct from its estimated ability and observed responses on the reduced subset using the performance-IRT (p-IRT) estimator (Polo et al., 2024) (detailed in Appendix F).
+P-IRT builds on the Test Characteristic Curve (Lord and Novick, 2008; Hambleton et al., 1991) which maps a model’s ability $\hat{\theta}$ to its expected probability of correctly answering benchmark items under the calibrated 3PL model, and refines this mapping by incorporating observed responses.
+As shown in Table 3, across all benchmarks, the reconstructed percent correct closely matches the raw full-bank percent correct, with MAE consistently below 5%, demonstrating that $\theta$ faithfully preserves global performance as measured by accuracy.
+This confirms that the latent ability estimates remain well aligned with conventional evaluation metrics.
+Moreover, ATLAS achieves this high reconstruction fidelity with substantially fewer administered items than competing methods.
+For example, on WinoGrande, ATLAS0.1 attains the lowest MAE (0.048) using only 70 items, compared to 133 items for MetaBench-Primary.
+4.3 Discriminating Models at Performance Extremes
+Figure 2 illustrates IRT’s key advantage: separating models with similar accuracy via latent ability estimates.
+Although accuracy and ability are highly correlated (0.99 on GSM8K; 0.96 on HellaSwag), systematic differences appear at performance extremes.
+In low-performing regimes, accuracy collapses into narrow bands (0.10–0.15), while IRT spans $\theta\approx-3$ to 1-1, distinguishing models that solve only easy items from those that handle harder ones.
+At the high end, ceiling effects compress accuracy, whereas IRT preserves discrimination across $\theta\approx 1.5$ to $2.5$.
+The right panels reveal substantial rank reordering: 23–31% of models shift by more than 10 positions when ranked by IRT rather than accuracy.
+Models that succeed on difficult items are ranked higher, while those relying on easy items are downweighted, yielding more reliable comparisons in saturated regimes.
+Similar trends hold across TruthfulQA, WinoGrande, and ARC (See Appendix G).
 4.4 Items Are Not Equally Informative
-Figure 3: Two models with identical accuracy (0.833) on WinoGrande receive different ability estimates ($\hat{\theta}_{A}=1.2$ vs $\hat{\theta}_{B}=0.6$).
-Model A succeeds on harder items (darker cells on right), while Model B answers easier items (darker cells on left).
-IRT captures these item difficulty patterns that raw accuracy cannot.
-Figure 3 demonstrates IRT’s key advantage: models with identical accuracy (0.833) receive different ability estimates ($\hat{\theta}_{A}=1.2$ vs $\hat{\theta}_{B}=0.6$) based on which items they solve correctly.
-Under the 3PL IRT model, items differ in discrimination $a_{i}$, difficulty $b_{i}$, and guessing $c_{i}$ parameters.
-Model A succeeds on high-difficulty items ($b_{i}>0.5$) with strong discrimination ($a_{i}>1.5$), yielding 2.3× more Fisher information than Model B, which mainly answers easier, less discriminative items ($b_{i}<-0.5$, $a_{i}<0.8$).
-IRT provides automatic quality control by weighting items according to their empirical contribution to distinguishing model abilities.
-In our calibrated banks, 3.2-5.7% of items exhibit negative discrimination ($a_{i}<0$), indicating systematic flaws where stronger models perform worse.
-For example, WinoGrande item #247 achieves 0.89 accuracy but $a_{i}=-0.43$ due to exploitable linguistic artifacts.
-Under raw accuracy, this flawed item contributes equally (weight = 1/N) to all scores, potentially inflating weak models.
-Under IRT, negative discrimination automatically down-weights its contribution by 82% (effective weight $\propto a_{i}^{2}\approx 0.18$), reducing measurement contamination and providing more reliable ability estimates than accuracy alone.
+Figure 3 illustrates a core principle of Item Response Theory: items contribute unequally to ability estimation.
+Although Models A and B achieve the same raw accuracy (0.833), their response patterns differ systematically across difficulty levels, leading to substantially different latent ability estimates ($\hat{\theta}_{A}=1.2$ vs.
+$\hat{\theta}_{B}=0.6$).
+Model A performs disproportionately well on harder items (D6–D9), which are more discriminative and thus carry greater informational value for estimating high ability, whereas Model B’s correct responses are concentrated on easier items (D1–D4), which provide limited evidence of advanced competence.
+As a result, identical accuracy masks meaningful differences in underlying capability: two models can answer the same number of questions correctly, yet differ substantially in latent ability because correct responses to difficult items provide strong evidence of high ability, while correct responses to easy items are expected and contribute little additional information.
+This example highlights why accuracy alone is an insufficient evaluation metric and how IRT leverages item difficulty to recover latent ability differences that raw accuracy fail to capture.
 Similar patterns of identical accuracy leading to different ability estimates are observed across ARC, HellaSwag, and other benchmarks, with additional heatmap visualizations provided in Appendix G.
 5 Discussion
-5.1 Implications for Benchmark Design
-Our findings underscore several critical considerations for the design of future large language model (LLM) benchmarks.
-First, reduced item sets necessitate content balancing to preserve assessment validity.
-Content balancing entails maintaining proportional representation across distinct skill domains or cognitive competencies to ensure that the benchmark reflects the intended construct rather than overemphasizing specific subskills.
-Without such balance, benchmarks risk producing biased or unrepresentative evaluations.
-The proposed adaptive framework can be readily extended to optimize both domain coverage and measurement precision simultaneously (cheng2009maximum, ).
-Achieving comparable balance using static, reduced item sets would be substantially more difficult and would require extensive manual tuning or domain expertise.
-Second, item quality exerts a direct influence on evaluation reliability.
-Our analysis reveals that 3.2–5.7% of items display negative discrimination parameters, indicating systematic flaws that undermine accuracy-based performance estimates.
-Although IRT automatically attenuates the influence of such items through probabilistic weighting, many existing benchmarks lack analogous quality control mechanisms.
-Consequently, poorly constructed or ambiguous items can distort aggregate scores and obscure genuine model differences.
-Future benchmark design should therefore adopt psychometric validation as standard practice, incorporating procedures such as discrimination screening and content alignment checks to enhance the interpretability and fairness of evaluation outcomes.
-Third, item exposure rate and test overlap rate are essential metrics for benchmark robustness and reusability.
-A low exposure rate ensures that individual items are rarely reused across assessments, reducing the risk of model memorization or contamination of test content.
-Similarly, a low test overlap rate—the proportion of shared items between distinct benchmark forms—preserves evaluation independence and minimizes redundancy.
-Adaptive benchmarking intrinsically achieves lower exposure and overlap rates by tailoring item selection to each model’s estimated proficiency, generating diverse yet psychometrically comparable test forms.
-This adaptivity confers a distinct advantage over static benchmarks, enabling more secure, scalable, and sustainable evaluation practices in iterative model development contexts.
-5.2 Limitations and Future Work
-Despite notable efficiency gains, our framework has several limitations.
-Initial calibration relies on a representative model population, which may become outdated as architectures evolve.
-Nonetheless, adaptive testing supports incremental updates: new model responses can be incorporated to refine item parameters and maintain calibration accuracy over time.
-The current implementation remains limited to multiple-choice formats, while generative or open-ended tasks require alternative scoring and modeling strategies.
-Moreover, our framework is unidimensional, assuming a single latent proficiency dimension across all items, whereas LLM capabilities are inherently multidimensional.
-Future work should extend the framework to multidimensional IRT formulations that jointly model reasoning, factuality, and linguistic ability.
-To facilitate broader adoption, we have implemented a modular scoring interface that allows users to define custom evaluation functions (e.g., determining whether a model answers a selected item correctly).
-Future work will explore online or Bayesian calibration techniques for continuous item updating, multidimensional modeling to capture diverse LLM capabilities, and hybrid adaptive–generative evaluation for open-ended tasks.
-We also plan to extend the system with interfaces for cross-model benchmarking, enhanced item exposure control, and visualization tools that improve interpretability and diagnostic insight.
+Our results underscore several psychometric principles that should guide the design of future LLM benchmarks.
+First, item quality is critical for reliable evaluation.
+Prior work has shown that mislabeled or ambiguous items are common in existing benchmarks (Vendrow et al., 2025; Gema et al., 2025).
+Although IRT naturally downweights such items through probabilistic modeling, static benchmarks lack mechanisms to prevent them from being sampled.
+Fluid Benchmarking (Hofmann et al., 2025) highlights this issue, showing that adaptive, information-based selection dramatically reduces the exposure of low-quality items.
+These findings collectively reinforce the need for psychometric validation in future benchmark design, including procedures such as discrimination screening and content-alignment checks.
+Second, model fit is essential for trustworthy item parameter estimation, yet it is rarely examined or reported in existing benchmark-reduction methods that apply IRT.
+When the underlying IRT model fits poorly, the resulting difficulty and discrimination estimates become unstable, which in turn compromises any downstream conclusions drawn from reduced item sets.
+Our results show that several widely used static subsets exhibit substantial misfit (Table 1), underscoring that model-fit diagnostics should be a standard requirement for any benchmark that applies IRT for item calibration or reduction.
+Routine reporting enables researchers to verify whether the assumed psychometric model adequately captures LLM response behavior before relying on the resulting item parameters or reduced test forms.
+Third, when large item banks are partitioned for computational efficiency, linking becomes necessary to prevent scale drift.
+Common-person linking places items on a shared latent scale, preserving interpretability and comparability.
+Such principled linking strategies are crucial for scalable benchmarking and for maintaining consistency across bank updates or domains.
+Finally, reduced benchmarks require explicit content balancing to maintain validity.
+Without constraints on domain coverage, reduced item sets may overemphasize narrow skills.
+Our adaptive framework naturally supports joint optimization of measurement precision and content balance (Cheng and Chang, 2009), whereas achieving similar guarantees in static subsets requires extensive manual tuning.
 6 Conclusion
-Adaptive testing replaces exhaustive benchmarking with efficient, contamination-resistant measurement.
-Our ATLAS framework reduces items by 90% while maintaining precision, separates models with identical accuracies using IRT-based ability estimates, and limits contamination through << 10% item exposure rates.
-With benchmark saturation and contamination risks increasing, adaptive testing provides a rigorous alternative, applying measurement science principles instead of static averaging.
+We presented ATLAS, an adaptive testing framework that reframes LLM evaluation by moving beyond static, accuracy-based benchmarking toward dynamic ability estimation.
+Through psychometric modeling and information-guided item selection, ATLAS achieves up to 90% item reduction, avoids accuracy ceiling effects, and reveals ability differences that static benchmarks overlook.
+Our analysis further highlights the importance of rigorous model-fit validation, item-quality assessment, and principled linking procedures for building reliable benchmarks, offering a more efficient, interpretable, and robust foundation to assess LLMs.
+Impact Statement
+This work aims to advance the methodology of large language model evaluation by introducing a psychometrically grounded, adaptive testing framework.
+By improving the reliability, efficiency, and interpretability of model assessment, our approach may support more informed benchmarking and comparison of LLMs.
+We do not foresee direct negative societal or ethical impacts arising from this work, as it focuses on evaluation rather than deployment or decision-making.
+Nonetheless, as with all benchmarking tools, care should be taken to ensure that evaluations are used responsibly and do not incentivize narrow optimization at the expense of broader model robustness.
 Appendix A Detailed Analysis of Average Score Limitations
@@
 Appendix B Comparison of IRT-Based Benchmark Methods
-Table 3: Comparison of IRT-Based Benchmark Methods for LLMs
+A detailed comparison of IRT-based benchmark methods is reported in Table 4.
 Factor TinyBenchmarks (Static) MetaBench (Static) ATLAS
@@
 Model Fit Rarely checked Rarely checked Possible fit checks
-Efficiency Fixed (100 items) Fixed (100 items) Fewer items, comparable accuracy
 Saturation Risk High High Low
-Appendix C Detailed Experimental Setup and Metrics
-C.1 Benchmarks and Datasets
+Table 4: Comparison of IRT-Based Benchmark Methods for LLMs
+Appendix C Data Preprocessing Details
+C.1 Point-Biserial Correlation Formula
+The point-biserial correlation (Allen and Yen, 2001) for item $i$ is defined as:
+$r_{pb}(i)\;=\;\frac{\bar{T}_{\,\ell\mid Y_{i\ell}=1}-\bar{T}_{\,\ell\mid Y_{i\ell}=0}}{s_{T}}\cdot\sqrt{p_{i}q_{i}},$
+where $\bar{T}_{\,\ell\mid Y_{i\ell}=1}$ and $\bar{T}_{\,\ell\mid Y_{i\ell}=0}$ are the mean total scores of models that answered item $i$ correctly and incorrectly, respectively; $s_{T}$ is the standard deviation of total scores $\{T_{\ell}\}$; $p_{i}=\tfrac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}Y_{i\ell}$ is the proportion of models that answered item $i$ correctly; $q_{i}=1-p_{i}$; and $|\mathcal{L}|$ is the number of models.
+C.2 Detailed Explanation of the Common-Person Calibration Procedure
+In this work, calibration refers to estimating item characteristics under the 3PL item response theory model.
+The purpose of calibration is to place all items on a shared difficulty scale so that performance comparisons across items and models become meaningful.
+Under the 3PL model, each item is described by a difficulty parameter, a discrimination parameter that captures how strongly the item differentiates between high- and low-performing models, and a guessing parameter that reflects the chance of a correct response when the model effectively guesses.
+When every LLM responds to the same items, their collective performance patterns allow these parameters to be estimated in a consistent way.
+This provides a principled way to identify which items are easy or difficult for models and which items are more or less informative.
+Calibrating the full benchmark at once would be computationally intensive because the 3PL model becomes more expensive to fit as the number of items grows.
+To make the process tractable, we divide the full item pool into several non-overlapping subsets and calibrate each subset independently.
+This reduces the computational load substantially, but it also means that each subset is estimated on its own internal scale.
+For example, the notion of “difficulty” in one subset is not automatically aligned with the notion of “difficulty” in another.
+A separate linking step is therefore required to place all subsets onto a shared scale.
+Linking requires shared reference points known as anchors.
+In educational measurement, anchors are typically common items or common examinees that appear in multiple test forms.
+They serve as a bridge that allows independently calibrated scales to be aligned.
+In our setting, every LLM responds to every item in the benchmark, which means that the same population of models appears in the calibration of each item subset.
+The models therefore act as common persons in the traditional psychometric sense and serve as the linking anchors for the benchmark.
+Their relative performance across subsets provides the information needed to align the scale of each subset with the others.
+The linking procedure examines how the same models perform across the different subsets and adjusts each subset’s scale so that the overall performance patterns match.
+If a model appears stronger than its peers in one subset and shows a similar relative standing in another, then the two subsets can be placed on the same scale by aligning the average performance level and the overall spread of performance.
+This rescaling is then applied to the item parameters of each subset so that all items, regardless of which subset they came from, are expressed on a single, unified difficulty metric.
+This form of common-person linking is particularly effective for LLM benchmarking.
+In human testing, it is rarely feasible for every examinee to respond to every item, and linking must rely on smaller or less reliable anchor sets.
+LLMs do not face constraints such as fatigue, practice effects, or time limits, which allows us to use the entire model population as a complete and stable set of anchors.
+This makes the linking process highly robust and enables a scalable calibration framework that achieves substantial computational efficiency while maintaining coherence across a very large item bank.
+WinoGrande TruthfulQA HellaSwag GSM8K ARC
+# Models used for calibration 4680 4635 3467 3775 3747
+# Models used for testing 521 516 386 420 417
+# Calibration subsets ($K$) 10 6 50 12 8
+# Items after filtering 1045 627 5600 1306 839
+Average RMSEA ($M_{2}$) 0.0565 0.0690 0.0482 0.0438 0.0595
+Table 5: Statistics describing the calibration dataset, testing dataset, item-bank size after filtering, number of calibration partitions (K), and average model–data fit (RMSEA from M2) across all benchmarks.
+C.3 Calibration Data, Item-Bank Partitioning, and Fit Statistics
+A summary of the calibration data, item-bank partitioning, and associated fit statistics is provided in Table 5.
+Appendix D Detailed Experimental Setup and Metrics
+D.1 Benchmarks and Datasets
 We conduct experiments on five diverse benchmarks covering different cognitive domains:
@@
 All experiments use the calibrated item banks from Section 3.3, ensuring consistent filtering and parameter quality across datasets.
-Items with poor psychometric properties (negative discrimination $a_{i}<0$, extreme difficulty $|b_{i}|>4$, or unrealistic guessing $c_{i}>0.5$) were excluded during calibration.
-C.2 Baseline Configurations
+D.2 Baseline Configurations
 We compare ATLAS against four fixed, non-adaptive strategies:
 Random Baseline: Samples 100 items uniformly from the full bank without consideration of item parameters or model ability.
-TinyBenchmarks: Uses the predetermined subset from [8], selected via clustering methods but without explicit Fisher information optimization for ability estimation.
-MetaBench-Primary and MetaBench-Secondary: Curated splits from [9] that require computationally expensive iterations to identify stable subsets.
+TinyBenchmarks: Uses the predetermined subset from Polo et al. (2024), selected via clustering methods but without explicit Fisher information optimization for ability estimation.
+MetaBench-Primary and MetaBench-Secondary: Curated splits from Kipnis et al.
+(2025) that require computationally expensive iterations to identify stable subsets.
 These splits emphasize predictive accuracy over psychometric validity.
@@
 Unlike ATLAS, these approaches do not adapt to individual test-takers and serve only as static reference points for accuracy–efficiency tradeoffs.
-C.3 ATLAS Configuration Details
+D.3 ATLAS Configuration Details
 For each model $\ell$, we run ATLAS under three precision-based stopping thresholds:
@@
 This setup balances precision and budget constraints, simulating realistic conditions for adaptive evaluation in production environments.
-C.4 Detailed Metric Definitions
-Average Mean Absolute Error (MAE):
-$\text{MAE}=\frac{1}{|\mathcal{M}|}\sum_{\ell\in\mathcal{M}}|\hat{\theta}_{\ell}-\hat{\theta}_{\ell}^{\text{whole}}|$
-where $\mathcal{M}$ is the set of evaluated models, $\hat{\theta}_{\ell}$ is the ATLAS-derived ability estimate, and $\hat{\theta}_{\ell}^{\text{whole}}$ is the reference estimate from the complete item bank.
-Average Item Exposure Rate:
-$\text{Exposure}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\frac{\text{count}(i)}{|\mathcal{M}|}$
-where $\text{count}(i)$ is the number of times item $i$ was administered across all models $\mathcal{M}$, and $|\mathcal{I}|$ is the total number of items in the bank.
-Test Overlap Rate:
-$\bar{Q}=\frac{1}{\binom{|\mathcal{M}|}{2}}\sum_{\ell_{1}<\ell_{2}}\frac{|\mathcal{T}_{\ell_{1}}\cap\mathcal{T}_{\ell_{2}}|}{\min(|\mathcal{T}_{\ell_{1}}|,|\mathcal{T}_{\ell_{2}}|)}$
-where $\mathcal{T}_{\ell}$ represents the set of items administered to model $\ell$.
-C.5 Psychometric Interpretation Guidelines
-Item Exposure Rate Benchmarks: Values below 20–30% are considered desirable in large-scale adaptive testing, as they indicate that no small subset of items dominates the evaluation, reducing security risks and promoting item bank longevity.
-Test Overlap Rate Benchmarks: Values under 10–15% are typically considered good in adaptive testing practice, providing sufficient form diversity to resist contamination while maintaining measurement consistency.
-Runtime Considerations: All measurements reflect computational efficiency in fully automated settings.
-Runtime includes item selection via Fisher information maximization but excludes model inference time, which varies significantly across architectures.
-Appendix D Model Fit Evaluation across Datasets
-Table 4 summarizes the average Root Mean Square Error of Approximation (RMSEA) from the MIRT M2 model fit across multiple benchmark datasets.
-Overall, all datasets exhibit low RMSEA values (ranging from 0.0436 to 0.0690), indicating a generally good fit between the latent multidimensional item response model and the observed response data.
-Among the evaluated datasets, GSM8K achieves the lowest average RMSEA (0.0436), suggesting that the underlying latent trait structure aligns particularly well with the response patterns, likely due to its consistent problem-solving format.
-In contrast, TruthfulQA shows a slightly higher RMSEA (0.0690), reflecting greater variability or multidimensionality in the response space.
-The narrow gaps between the 5% and 95% RMSEA bounds across datasets indicate model stability and relatively low dispersion of fit indices among individual files.
-This consistency across diverse benchmarks such as Winogrande, HellaSwag, and ARC supports the reliability of the MIRT calibration process.
-Taken together, these results suggest that the applied MIRT M2 model provides a satisfactory and robust representation of item-level relationships across heterogeneous reasoning and language understanding tasks.
-Table 4: Summary of MIRT M2 Average RMSEA across datasets
-Dataset Avg. RMSEA Avg. 5% RMSEA Avg. 95% RMSEA Num. Files
-Winogrande $0.0565$ $0.0561$ $0.0568$ $10$
-TruthfulQA $0.0690$ $0.0686$ $0.0693$ 66
-HellaSwag $0.0482$ $0.0478$ $0.0486$ $50$
-GSM8K $0.0436$ $0.0433$ $0.0440$ $12$
-ARC $0.0588$ $0.0585$ $0.0592$ 88
-  • Note: ‘Avg. RMSEA‘ is the mean RMSEA across all files for each dataset. ‘Avg. 5% RMSEA‘ and ‘Avg. 95% RMSEA‘ are the averages of the file-level 5th and 95th percentile RMSEA bounds, respectively.
-Appendix E Data Preprocessing Details
-E.1 Point-Biserial Correlation Formula
-The point-biserial correlation [23] for item $i$ is defined as:
-$r_{pb}(i)\;=\;\frac{\bar{T}_{\,\ell\mid Y_{i\ell}=1}-\bar{T}_{\,\ell\mid Y_{i\ell}=0}}{s_{T}}\cdot\sqrt{p_{i}q_{i}},$
-where $\bar{T}_{\,\ell\mid Y_{i\ell}=1}$ and $\bar{T}_{\,\ell\mid Y_{i\ell}=0}$ are the mean total scores of models that answered item $i$ correctly and incorrectly, respectively; $s_{T}$ is the standard deviation of total scores $\{T_{\ell}\}$; $p_{i}=\tfrac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}Y_{i\ell}$ is the proportion of models that answered item $i$ correctly; $q_{i}=1-p_{i}$; and $|\mathcal{L}|$ is the number of models.
-E.2 Filtering Results
-WinoGrande TruthfulQA HellaSwag GSM8K ARC
-Num. Selected Models 4680 4635 3467 4195 4162
-Num. Selected Item 1046 628 5608 1306 842
-Avg. $M_{2}$ RMSEA 0.0565 0.0690 0.0482 0.0436 0.0588
-Table 5: Summary of model and item filtering per benchmark.
-RMSEA values reflect average fit from $M_{2}$ statistics under the 3PL IRT model.
-Calibration and ability estimation were performed using the mirt package in R, with WLE as the estimator.
-Items were filtered based on variability, ceiling effects, and point-biserial discrimination.
-Appendix F Evaluation Metric Definitions
+Appendix E Evaluation Metric Definitions
 This section provides the exact mathematical definitions of the evaluation metrics introduced in Section 4, along with brief interpretations.
-Average Mean Absolute Error (MAE).
-For each model $\ell$, let $\hat{\theta}_{\ell}$ denote the CAT-derived ability estimate and $\hat{\theta}_{\ell}^{\text{whole}}$ the full-bank reference estimate. The average MAE is:
-$\text{MAE}=\frac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}\left|\hat{\theta}_{\ell}-\hat{\theta}_{\ell}^{\text{whole}}\right|,$ (4)
-where $|\mathcal{L}|$ is the number of models evaluated. Interpretation: Lower MAE indicates higher fidelity, i.e., CAT ability estimates more closely match the whole-bank reference.
+Average Mean Absolute Error (MAE) and Standard Error (SE).
+We compute MAE for both ability estimates and accuracy scores.
+For ability, let $\hat{\theta}_{\ell}$ denote the CAT-derived estimate and $\hat{\theta}_{\ell}^{\mathrm{whole}}$ the full-bank reference.
+The ability MAE is
+$\text{MAE}_{\theta}=\frac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}\left|\hat{\theta}_{\ell}-\hat{\theta}_{\ell}^{\mathrm{whole}}\right|.$
+For accuracy, let $\widehat{Acc}^{\mathrm{p\text{-}IRT}}_{\ell}$ denote the reconstructed accuracy (e.g., via p-IRT) and $\mathrm{Acc}_{\ell}^{\mathrm{raw}}$ the observed raw accuracy.
+The accuracy MAE is
+$\text{MAE}_{\mathrm{acc}}=\frac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}\left|\widehat{Acc}^{\mathrm{p\text{-}IRT}}_{\ell}-\mathrm{Acc}_{\ell}^{\mathrm{raw}}\right|.$
+To quantify variability across models, we also report the standard error (SE) of MAE. Let $e_{\ell}$ denote the per-model absolute error and $\overline{e}$ its mean. The standard deviation (SD) is
+$\text{SD}=\sqrt{\frac{1}{|\mathcal{L}|-1}\sum_{\ell\in\mathcal{L}}(e_{\ell}-\overline{e})^{2}},$
+and the standard error (SE) is
+$\text{SE}=\frac{\text{SD}}{\sqrt{|\mathcal{L}|}}.$
+Interpretation: Lower MAE and SE indicate higher fidelity and greater stability: CAT-derived estimates more closely match whole-bank references (for ability) or observed scores (for accuracy) and do so consistently across models.
+Information Efficiency Score (IES).
+To compare the efficiency of different evaluation methods, we define the Information Efficiency Score (IES) relative to a baseline of 100-item uniform random sampling (Random_100).
+For a given method, let $\text{MAE}_{\text{method}}$ denote its average MAE and $\text{MAE}_{\text{Random}}$ the MAE under Random_100.
+Let $\text{Items}_{\text{method}}$ denote the average number of selected subset items.
+The IES is:
+$\text{IES}\;=\;\left(\frac{\text{MAE}_{\text{method}}}{\text{MAE}_{\text{Random}}}\right)\left(\frac{\text{Items}_{\text{method}}}{100}\right).$ (4)
+Interpretation: An IES value below 1 indicates that the method achieves a better accuracy–efficiency tradeoff than the Random_100 baseline, requiring fewer items and/or producing lower error for the same number of items.
+An IES value of 1 means the method is equally efficient as Random_100.
+Values greater than 1 indicate lower efficiency, meaning the method uses more items and/or yields higher error than the baseline.
 Average Item Exposure Rate.
@@
 Test Overlap Rate.
-Following chen2005simplified [24], the expected proportion of common items between two randomly selected test forms is given by
+Following Hsu et al. (2013), the expected proportion of common items between two randomly selected test forms is given by
 $\bar{Q}=\frac{|\mathcal{L}|\sum_{i=1}^{|\mathcal{I}|}P(A_{i})^{2}}{\bar{L}(|\mathcal{L}|-1)}-\frac{1}{|\mathcal{L}|-1},$ (7)
-where $\bar{L}$ is the average test length. Interpretation: Lower values of $\bar{Q}$ imply greater test form diversity, which reduces risks of collusion and item memorization.
-Correlation Metrics
+where $\bar{L}$ is the average test length.
+Interpretation: Lower values of $\bar{Q}$ imply greater test form diversity, which reduces risks of collusion and item memorization.
+E.1 Correlation Metrics
 For completeness, we provide the definitions of the rank-based correlation coefficients used in Section 4.
@@
 $\tau=\frac{(\#\text{concordant pairs})-(\#\text{discordant pairs})}{\tfrac{1}{2}n(n-1)}.$
+Appendix F Performance-IRT (p-IRT) Estimator
+The Performance-IRT (p-IRT) estimator (Polo et al., 2024) is a probabilistic scoring method used to compute expected accuracy when only a subset of benchmark items is observed.
+Conceptually, p-IRT is grounded in the Test Characteristic Curve (TCC) (Lord and Novick, 2008; Hambleton et al., 1991), which maps a model’s ability $\hat{\theta}$ to its expected probability of correctly answering items under the calibrated 3PL model.
+It provides an estimate of a model’s overall benchmark accuracy without requiring evaluation on the full item set.
+Goal.
+We formulate LLM evaluation as a psychometric measurement problem.
+Let $\mathcal{I}$ denote the full set of benchmark items and $\mathcal{L}$ the set of language models.
+For each model $\ell\in\mathcal{L}$ and item $i\in\mathcal{I}$, we observe a binary response $Y_{i,\ell}\in\{0,1\}$, forming the item–response matrix $\{Y_{i,\ell}\}_{i\in\mathcal{I},\,\ell\in\mathcal{L}}$.
+The true full-benchmark accuracy of model $\ell$ is
+$\mathrm{Acc}_{\ell}^{\mathrm{raw}}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}Y_{i,\ell}.$
+The p-IRT estimator approximates $\mathrm{Acc}_{\ell}^{\mathrm{raw}}$ when only a subset of items $\widehat{\mathcal{I}}\subseteq\mathcal{I}$ is observed, by combining the model’s observed responses on $\widehat{\mathcal{I}}$ with IRT-predicted probabilities on the remaining items $\mathcal{I}\setminus\widehat{\mathcal{I}}$.
+Estimator.
+P-IRT computes the conditional expectation
+$\widehat{Acc}^{\mathrm{p\text{-}IRT}}_{\ell}=\mathbb{E}\!\left[\mathrm{Acc}_{\ell}^{\mathrm{raw}}\,\middle|\,\{Y_{i,\ell}:i\in\widehat{\mathcal{I}}\}\right$],
+which is the minimum–mean-squared-error predictor of $\mathrm{Acc}_{\ell}^{\mathrm{raw}}$ under the calibrated IRT model. Then
+$\widehat{Acc}^{\mathrm{p\text{-}IRT}}_{\ell}=\frac{|\widehat{\mathcal{I}}|}{|\mathcal{I}|}\cdot\underbrace{\frac{1}{|\widehat{\mathcal{I}}|}\sum_{i\in\widehat{\mathcal{I}}}Y_{i,\ell}}_{\text{Observed accuracy}}\;+\;\frac{|\mathcal{I}\setminus\widehat{\mathcal{I}}|}{|\mathcal{I}|}\cdot\underbrace{\frac{1}{|\mathcal{I}\setminus\widehat{\mathcal{I}}|}\sum_{i\in\mathcal{I}\setminus\widehat{\mathcal{I}}}\hat{p}_{i,\ell}}_{\text{Unobserved TCC}},$
+where
+$\hat{p}_{i,\ell}=P\!\left(Y_{i,\ell}=1\mid\hat{\theta}_{\ell},\hat{a}_{i},\hat{b}_{i},\hat{c}_{i}\right)$
+is the predicted probability of correctness for model $\ell$ under the calibrated 3PL model.
+Intuition.
+The p-IRT estimator is a weighted combination of:
+  • Observed accuracy on the subset $\widehat{\mathcal{I}}$.
+  • Unobserved TCC on the remaining items $\mathcal{I}\setminus\widehat{\mathcal{I}}$, based on the model’s ability $\hat{\theta}_{\ell}$ and item parameters.
+The weight $|\widehat{\mathcal{I}}|/|\mathcal{I}|\in[0,1]$ corresponds to the proportion of observed items and determines the tradeoff between observed and predicted performance.
+Use in This Work.
+We apply p-IRT to reconstruct accuracy from ability estimates $\hat{\theta}_{\ell}$.
+As shown in Table 3, reconstructed accuracies closely match raw accuracies across benchmarks, confirming that ability estimates retain the global performance structure while smoothing noise and offering finer discrimination than raw accuracy alone.
 Appendix G Additional Experimental Results
-While both TinyBenchmarks [8] and MetaBench [9] include MMLU [25] as part of their evaluation suites, they treat it as a single unified dataset by aggregating all 57 subject areas.
+G.1 Additional Benchmark Results
+Tables 6 and 7 present results for GSM8K and ARC benchmarks, complementing the main results in Tables 2 and 3.
+Method GSM8K ARC
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Method GSM8K ARC
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+G.2 Interpreting Test Overlap, Exposure, and Runtime in ATLAS
+Table 8 provides a detailed breakdown of adaptive evaluation behavior across all five benchmarks, summarizing test overlap, average item exposure, and selection time.
+These metrics together illustrate how ATLAS balances efficiency, diversity, and computational scalability when administering adaptive tests.
+Formal definitions of all metrics are included in Appendix E.
+Test overlap rates quantify how frequently different models are exposed to the same items.
+Across benchmarks, overlap remains modest, ranging from roughly 11% to 23%.
+These values are far lower than those of static subsets, which administer identical items to all models.
+The relatively low overlap indicates that ATLAS tailors item sequences to each model’s evolving ability estimate rather than relying on a fixed set of questions.
+For example, HellaSwag reaches the lowest overlap values (as low as 11.26%), reflecting its large item pool and the wide range of informative items available.
+Higher overlap on datasets such as GSM8K (approximately 20–24%) reflects the smaller bank size and the concentration of discriminative items in particular ability regions.
+Overall, the overlap statistics confirm that ATLAS provides genuine adaptivity while preserving comparability across models.
+Average item exposure rates remain low across all settings, consistently under 12% and often much lower.
+Exposure values around 3–5% on HellaSwag and WinoGrande (for SE thresholds of 0.2 and 0.3) indicate that ATLAS does not rely excessively on a small subset of items.
+Low exposure reduces the risk of memorization or contamination in long-term evaluation scenarios, broadens the portion of the item bank that contributes to measurement, and ensures that no individual item disproportionately influences ability estimation.
+The pattern across SE thresholds reflects a standard property of adaptive testing: when fewer items are required (larger SE thresholds), exposure becomes more concentrated on the most informative items.
+In ATLAS, this concentration remains moderate, indicating healthy rotation among informative items.
+Selection time reflects the computational cost of the full adaptive selection loop, including Fisher information computation and termination checks.
+It refers to the complete runtime of the adaptive item selection loop for each model, rather than the time required for a single item decision.
+Times range from 9 to 76 seconds per model and scale predictably with benchmark size.
+TruthfulQA, with 628 items, achieves the fastest selection times (approximately 9–16 seconds).
+In contrast, HellaSwag, with more than 5600 items, shows the longest selection times (57–76 seconds), due to the larger number of items evaluated when determining the most informative question at each step.
+Importantly, even in the largest setting, selection remains well under 90 seconds, and for all other benchmarks it typically completes within tens of seconds.
+This confirms that ATLAS is computationally practical for both interactive evaluation and large-scale benchmarking workflows.
+Taken together, the results in Table 8 show that ATLAS achieves high adaptivity, broad item utilization, and practical runtime efficiency across diverse benchmarks.
+Low overlap and exposure promote content coverage and robustness, while stable runtime performance ensures operational scalability without compromising statistical quality.
+Benchmark Method Test Overlap $\downarrow$ Avg. Item $\downarrow$ Avg. Selection $\downarrow$
+(Item #) Rate (%) Exposure Rate (%) Time (s)
+WinoGrande (1046) ATLASSE≤0.1 18.22 8.24 40.99
+ATLASSE≤0.2 14.93 4.71 19.92
+ATLASSE≤0.3 17.03 4.04 16.74
+TruthfulQA (628) ATLASSE≤0.1 17.32 7.86 15.97
+ATLASSE≤0.2 18.43 9.58 9.37
+ATLASSE≤0.3 18.07 9.49 9.72
+HellaSwag (5600) ATLASSE≤0.1 11.26 3.86 75.52
+ATLASSE≤0.2 13.72 4.78 56.93
+ATLASSE≤0.3 13.67 4.82 57.06
+GSM8K (1307) ATLASSE≤0.1 21.27 7.21 45.69
+ATLASSE≤0.2 20.78 4.40 24.08
+ATLASSE≤0.3 23.70 5.54 19.06
+ARC (842) ATLASSE≤0.1 19.15 11.15 30.99
+ATLASSE≤0.2 17.09 5.41 13.98
+ATLASSE≤0.3 19.60 9.21 11.82
+Table 8: Adaptive evaluation efficiency and diversity.
+ATLAS maintains low item exposure rates ($12\%$) and moderate test overlap ($13-24\%$) with fast selection times (<76<76 seconds per model).
+Lower values indicate better performance for all metrics.
+G.3 Experiment on MMLU
+While both TinyBenchmarks (Polo et al., 2024) and MetaBench (Kipnis et al., 2025) include MMLU (Hendrycks et al., 2020) as part of their evaluation suites, they treat it as a single unified dataset by aggregating all 57 subject areas.
 In contrast, we perform evaluation on a per-subject basis.
@@
 Aggregating across all subjects can obscure these domain-specific patterns and limit interpretability in adaptive assessment.
-Since each MMLU subject contains approximately 150 items, we randomly selected several representative subjects—Anatomy, Astronomy, and High-School Macroeconomics—to illustrate ATLAS’s behavior under data-sparse conditions.
-The corresponding results are reported in Table 6.
+The corresponding results are reported in Table LABEL:tab:mmlu_adaptive_results.
 Despite the small number of items per subject, ATLAS consistently demonstrates robust adaptive evaluation performance.
@@
 Benchmark Method MAE $\downarrow$ Avg. Item $\downarrow$ Test Overlap $\downarrow$ Exposure Rate $\downarrow$ Avg. Time (s) $\downarrow$
-MMLU-Anatomy
-(113 items)
-ATLASSE≤0.1 0.099 100.00 0.94 0.88 4.93
+MMLU-Abstract Algebra (79 items) ATLASSE≤0.1 0.025 52.45 0.6808 0.6639 0.59
+ATLASSE≤0.2 0.067 32.18 0.4413 0.4073 0.36
+ATLASSE≤0.3 0.098 19.90 0.3169 0.2519 0.22
+MMLU-Anatomy (113 items) ATLASSE≤0.1 0.099 100.00 0.94 0.88 4.93
 ATLASSE≤0.2 0.149 53.29 0.54 0.47 2.70
 ATLASSE≤0.3 0.235 20.49 0.32 0.18 0.98
-MMLU-Astronomy
-(136 items)
-ATLASSE≤0.1 0.099 93.22 0.82 0.69 6.30
+MMLU-Astronomy (136 items) ATLASSE≤0.1 0.099 93.22 0.82 0.69 6.30
 ATLASSE≤0.2 0.157 48.21 0.48 0.35 3.29
 ATLASSE≤0.3 0.235 11.80 0.40 0.11 0.79
-MMLU-High School
-Macroeconomics (281 items)
-ATLASSE≤0.1 0.089 87.94 0.48 0.31 10.49
+MMLU-Business Ethics (95 items) ATLASSE≤0.1 0.040 86.61 0.9122 0.9117 4.09
+ATLASSE≤0.2 0.072 54.43 0.5912 0.5729 5.29
+ATLASSE≤0.3 0.160 13.46 0.4361 0.1417 0.63
+MMLU-Clinical Knowledge (198 items) ATLASSE≤0.1 0.044 99.86 0.7302 0.5043 8.69
+ATLASSE≤0.2 0.186 23.33 0.2894 0.1477 4.20
+ATLASSE≤0.3 0.244 10.46 0.2537 0.1016 0.91
+MMLU-College Biology (131 items) ATLASSE≤0.1 0.031 100.00 0.8870 0.7634 5.98
+ATLASSE≤0.2 0.083 67.87 0.6168 0.5186 4.20
+ATLASSE≤0.3 0.134 29.40 0.3380 0.2315 1.90
+MMLU-College Chemistry (77 items) ATLASSE≤0.1 0.029 77.00 1.0000 1.0000 0.72
+ATLASSE≤0.2 0.070 52.23 0.6857 0.6783 0.47
+ATLASSE≤0.3 0.116 17.84 0.4230 0.2308 0.25
+MMLU-College Computer Science (84 items) ATLASSE≤0.1 0.030 82.71 0.9845 0.9844 0.73
+ATLASSE≤0.2 0.066 41.67 0.5305 0.4961 0.39
+ATLASSE≤0.3 0.094 13.59 0.3367 0.1621 0.19
+MMLU-College Mathematics (69 items) ATLASSE≤0.1 0.025 63.86 0.9267 0.9256 0.62
+ATLASSE≤0.2 0.062 44.67 0.6669 0.6478 0.43
+ATLASSE≤0.3 0.089 28.48 0.4832 0.4128 0.29
+MMLU-College Medicine (157 items) ATLASSE≤0.1 0.038 100.00 0.8039 0.6369 7.02
+ATLASSE≤0.2 0.111 66.64 0.5986 0.4350 4.90
+ATLASSE≤0.3 0.171 27.77 0.3062 0.2061 2.02
+MMLU-College Physics (72 items) ATLASSE≤0.1 0.020 69.39 0.9634 0.9632 0.69
+ATLASSE≤0.2 0.041 61.78 0.8618 0.8580 0.63
+ATLASSE≤0.3 0.074 30.69 0.5167 0.4264 0.36
+MMLU-Computer Security (84 items) ATLASSE≤0.1 0.019 84.00 1.0000 1.0000 0.71
+ATLASSE≤0.2 0.034 75.99 0.9067 0.9049 0.65
+ATLASSE≤0.3 0.065 29.54 0.4484 0.3508 0.35
+MMLU-Conceptual Physics (203 items) ATLASSE≤0.1 0.042 69.43 0.5217 0.3417 4.02
+ATLASSE≤0.2 0.160 13.86 0.1844 0.0713 0.98
+ATLASSE≤0.3 0.205 10.58 0.1844 0.0596 0.78
+MMLU-Econometrics (102 items) ATLASSE≤0.1 0.031 91.67 0.9026 0.8992 5.38
+ATLASSE≤0.2 0.074 56.23 0.5678 0.5508 4.05
+ATLASSE≤0.3 0.148 13.44 0.4002 0.1442 0.64
+MMLU-Electrical Engineering (126 items) ATLASSE≤0.1 0.032 100.00 0.8917 0.7937 5.98
+ATLASSE≤0.2 0.081 61.94 0.5710 0.4917 4.05
+ATLASSE≤0.3 0.134 18.70 0.3021 0.1498 1.05
+MMLU-Elementary Mathematics (220 items) ATLASSE≤0.1 0.050 66.10 0.4546 0.3006 3.80
+ATLASSE≤0.2 0.191 13.34 0.2313 0.0843 0.87
+ATLASSE≤0.3 0.237 10.02 0.2020 0.0652 0.67
+MMLU-Formal Logic (109 items) ATLASSE≤0.1 0.030 91.05 0.8751 0.8359 5.43
+ATLASSE≤0.2 0.098 31.64 0.3807 0.2895 2.18
+ATLASSE≤0.3 0.146 20.89 0.2762 0.1919 1.26
+MMLU-Global Facts (81 items) ATLASSE≤0.1 0.022 65.09 0.8060 0.8041 0.56
+ATLASSE≤0.2 0.038 57.62 0.7158 0.7115 0.50
+ATLASSE≤0.3 0.069 18.03 0.4382 0.2225 0.22
+MMLU-High School Biology (251 items) ATLASSE≤0.1 0.037 100.00 0.6336 0.3984 6.09
+ATLASSE≤0.2 0.124 37.74 0.2869 0.1932 2.78
+ATLASSE≤0.3 0.171 25.49 0.2193 0.1480 1.98
+MMLU-High School Chemistry (169 items) ATLASSE≤0.1 0.033 100.00 0.7729 0.5917 5.64
+ATLASSE≤0.2 0.111 36.56 0.3257 0.2349 2.62
+ATLASSE≤0.3 0.151 17.30 0.2078 0.1182 1.00
+MMLU-High School Computer Science (94 items) ATLASSE≤0.1 0.021 94.00 1.0000 1.0000 0.74
+ATLASSE≤0.2 0.041 73.23 0.7911 0.7788 0.59
+ATLASSE≤0.3 0.066 31.34 0.4380 0.3330 0.34
+MMLU-High School European History (147 items) ATLASSE≤0.1 0.038 100.00 0.8630 0.6803 6.31
+ATLASSE≤0.2 0.081 75.10 0.6755 0.5368 5.16
+ATLASSE≤0.3 0.137 27.05 0.3238 0.2001 2.15
+MMLU-High School Geography (170 items) ATLASSE≤0.1 0.036 85.66 0.6861 0.5039 4.89
+ATLASSE≤0.2 0.095 45.02 0.4002 0.2904 2.90
+ATLASSE≤0.3 0.142 17.80 0.3141 0.1356 1.19
+MMLU-High School Government & Politics (154 items) ATLASSE≤0.1 0.033 100.00 0.8509 0.6494 5.94
+ATLASSE≤0.2 0.062 92.63 0.7887 0.6011 5.54
+ATLASSE≤0.3 0.124 26.66 0.3425 0.1989 1.69
+MMLU-High School Mathematics (199 items) ATLASSE≤0.1 0.044 56.19 0.4137 0.2822 3.31
+ATLASSE≤0.2 0.094 47.18 0.3732 0.2486 2.86
+ATLASSE≤0.3 0.183 12.12 0.2503 0.0894 0.79
+MMLU-High School Microeconomics (193 items) ATLASSE≤0.1 0.040 93.17 0.7021 0.4828 5.55
+ATLASSE≤0.2 0.174 18.27 0.2907 0.1188 1.25
+ATLASSE≤0.3 0.223 10.64 0.2579 0.1005 0.79
+MMLU-High School Macroeconomics (281 items) ATLASSE≤0.1 0.089 87.94 0.48 0.31 10.49
 ATLASSE≤0.2 0.197 27.64 0.22 0.13 3.46
 ATLASSE≤0.3 0.240 18.26 0.20 0.09 2.34
+MMLU-High School Physics (119 items) ATLASSE≤0.1 0.026 63.30 0.5825 0.5317 3.54
+ATLASSE≤0.2 0.055 51.75 0.4900 0.4356 3.08
+ATLASSE≤0.3 0.106 20.30 0.3393 0.1932 1.37
+MMLU-High School Psychology (285 items) ATLASSE≤0.1 0.035 100.00 0.6086 0.3509 6.19
+ATLASSE≤0.2 0.147 26.80 0.2433 0.1284 1.97
+ATLASSE≤0.3 0.196 11.48 0.2101 0.0881 0.92
+MMLU-High School Statistics (185 items) ATLASSE≤0.1 0.038 69.47 0.5587 0.3757 4.07
+ATLASSE≤0.2 0.154 21.48 0.2987 0.1219 1.67
+ATLASSE≤0.3 0.205 11.81 0.2483 0.0700 1.03
+MMLU-High School US History (183 items) ATLASSE≤0.1 0.030 96.68 0.7761 0.5283 5.77
+ATLASSE≤0.2 0.069 73.31 0.6137 0.4027 4.80
+ATLASSE≤0.3 0.151 13.62 0.3348 0.1319 1.05
+MMLU-High School World History (206 items) ATLASSE≤0.1 0.031 96.43 0.7407 0.4680 5.88
+ATLASSE≤0.2 0.108 40.83 0.3806 0.2580 2.85
+ATLASSE≤0.3 0.176 17.10 0.2629 0.1230 1.28
+MMLU-Human Aging (188 items) ATLASSE≤0.1 0.034 100.00 0.7341 0.5319 5.86
+ATLASSE≤0.2 0.108 40.24 0.3546 0.2515 2.69
+ATLASSE≤0.3 0.154 27.38 0.2445 0.1802 2.00
+MMLU-Human Sexuality (116 items) ATLASSE≤0.1 0.032 100.00 0.9137 0.8621 6.07
+ATLASSE≤0.2 0.062 83.66 0.7655 0.7212 5.25
+ATLASSE≤0.3 0.154 20.43 0.2825 0.1759 1.51
+MMLU-International Law (103 items) ATLASSE≤0.1 0.028 100.00 0.9780 0.9709 5.88
+ATLASSE≤0.2 0.061 76.17 0.7582 0.7400 4.72
+ATLASSE≤0.3 0.105 34.28 0.4298 0.3333 2.57
+MMLU-Jurisprudence (99 items) ATLASSE≤0.1 0.021 99.00 1.0000 1.0000 0.79
+ATLASSE≤0.2 0.047 71.20 0.7262 0.7191 0.60
+ATLASSE≤0.3 0.095 33.57 0.3986 0.3399 0.38
+MMLU-Logical Fallacies (147 items) ATLASSE≤0.1 0.030 100.00 0.8562 0.6803 5.92
+ATLASSE≤0.2 0.064 70.48 0.6483 0.4928 4.73
+ATLASSE≤0.3 0.117 22.21 0.2938 0.1632 1.69
+MMLU-Machine Learning (98 items) ATLASSE≤0.1 0.027 74.29 0.7619 0.7586 3.34
+ATLASSE≤0.2 0.069 39.44 0.4402 0.4018 1.89
+ATLASSE≤0.3 0.143 13.81 0.3032 0.1410 0.71
+MMLU-Management (92 items) ATLASSE≤0.1 0.022 92.00 1.0000 1.0000 0.74
+ATLASSE≤0.2 0.024 90.39 0.9831 0.9831 0.72
+ATLASSE≤0.3 0.059 44.91 0.5238 0.4881 0.47
+MMLU-Marketing (178 items) ATLASSE≤0.1 0.033 100.00 0.7974 0.5618 5.78
+ATLASSE≤0.2 0.081 61.18 0.5305 0.3645 4.06
+ATLASSE≤0.3 0.134 25.09 0.3035 0.1660 1.84
+MMLU-Medical Genetics (87 items) ATLASSE≤0.1 0.024 84.38 0.9699 0.9698 0.74
+ATLASSE≤0.2 0.045 69.31 0.7995 0.7967 0.62
+ATLASSE≤0.3 0.072 55.16 0.6449 0.6349 0.51
+MMLU-Miscellaneous (314 items) ATLASSE≤0.1 0.048 65.51 0.3154 0.2085 3.66
+ATLASSE≤0.2 0.156 24.00 0.1576 0.0991 1.58
+ATLASSE≤0.3 0.203 15.73 0.1295 0.0697 1.10
+MMLU-Moral Disputes (229 items) ATLASSE≤0.1 0.041 99.73 0.6574 0.4355 6.03
+ATLASSE≤0.2 0.142 27.23 0.2498 0.1526 1.94
+ATLASSE≤0.3 0.195 13.45 0.2077 0.1090 0.98
+MMLU-Moral Scenarios (486 items) ATLASSE≤0.1 0.059 34.18 0.1610 0.0704 2.24
+ATLASSE≤0.2 0.114 19.23 0.1216 0.0436 1.41
+ATLASSE≤0.3 0.158 14.14 0.0986 0.0334 1.05
+MMLU-Nutrition (244 items) ATLASSE≤0.1 0.039 95.50 0.6325 0.3914 6.00
+ATLASSE≤0.2 0.129 38.88 0.3028 0.1953 2.89
+ATLASSE≤0.3 0.176 21.25 0.2221 0.1191 1.69
+MMLU-Philosophy (214 items) ATLASSE≤0.1 0.041 96.45 0.6360 0.4510 5.90
+ATLASSE≤0.2 0.152 23.09 0.1983 0.1258 1.89
+ATLASSE≤0.3 0.202 16.56 0.1657 0.0927 1.37
+MMLU-Prehistory (265 items) ATLASSE≤0.1 0.039 93.64 0.6004 0.3531 5.72
+ATLASSE≤0.2 0.099 44.74 0.4126 0.2078 3.28
+ATLASSE≤0.3 0.190 12.29 0.2773 0.0989 0.95
+MMLU-Professional Accounting (221 items) ATLASSE≤0.1 0.047 63.18 0.4458 0.2861 3.83
+ATLASSE≤0.2 0.170 14.16 0.2016 0.0682 1.03
+ATLASSE≤0.3 0.220 10.46 0.1899 0.0647 0.83
+MMLU-Professional Law (518 items) ATLASSE≤0.1 0.064 26.94 0.1811 0.0519 1.84
+ATLASSE≤0.2 0.151 10.88 0.1380 0.0283 0.98
+ATLASSE≤0.3 0.188 10.06 0.1308 0.0271 0.90
+MMLU-Professional Medicine (218 items) ATLASSE≤0.1 0.044 95.14 0.6919 0.4363 6.11
+ATLASSE≤0.2 0.167 24.91 0.3403 0.1649 2.12
+ATLASSE≤0.3 0.220 10.25 0.2794 0.1019 0.95
+MMLU-Professional Psychology (344 items) ATLASSE≤0.1 0.051 58.00 0.3224 0.1685 3.86
+ATLASSE≤0.2 0.166 11.53 0.1850 0.0670 1.06
+ATLASSE≤0.3 0.205 10.08 0.1715 0.0645 0.95
+MMLU-Public Relations (93 items) ATLASSE≤0.1 0.027 93.05 1.0000 1.0000 0.74
+ATLASSE≤0.2 0.048 72.97 0.7926 0.7854 0.61
+ATLASSE≤0.3 0.095 24.50 0.3729 0.2633 0.31
+MMLU-Security Studies (201 items) ATLASSE≤0.1 0.038 100.00 0.7247 0.4975 6.11
+ATLASSE≤0.2 0.119 34.74 0.3103 0.2038 2.59
+ATLASSE≤0.3 0.171 17.12 0.1992 0.1103 1.30
+MMLU-Sociology (166 items) ATLASSE≤0.1 0.032 100.00 0.8412 0.6024 6.02
+ATLASSE≤0.2 0.070 72.92 0.6391 0.5095 5.00
+ATLASSE≤0.3 0.137 37.22 0.3946 0.2712 2.70
+MMLU-US Foreign Policy (90 items) ATLASSE≤0.1 0.024 90.00 1.0000 1.0000 0.72
+ATLASSE≤0.2 0.026 89.55 0.9950 0.9950 0.71
+ATLASSE≤0.3 0.063 30.48 0.4204 0.3389 0.37
+MMLU-Virology (140 items) ATLASSE≤0.1 0.032 100.00 0.8644 0.7143 6.01
+ATLASSE≤0.2 0.081 63.64 0.5787 0.4711 4.40
+ATLASSE≤0.3 0.134 24.11 0.2965 0.1926 1.84
+MMLU-World Religions (137 items) ATLASSE≤0.1 0.033 100.00 0.8796 0.7299 6.05
+ATLASSE≤0.2 0.068 87.19 0.7734 0.6657 5.23
+ATLASSE≤0.3 0.134 37.56 0.4063 0.2984 2.68
+TruthfulQA
 Figure 4: Comparison of raw average scores and whole-bank ability estimates on TruthfulQA.
@@
 (Right) Rank comparison shows strong consistency between the two measures (Spearman $\rho=0.97$, Kendall $\tau=0.87$), but ability-based ranking provides finer resolution, especially in distinguishing weaker and stronger models beyond what raw accuracy captures.
+WinoGrande
 Figure 5: Comparison of raw average scores and whole-bank ability estimates on WinoGrande.
-(Left) Whole-bank estimates show a non-linear relationship with average score and reveal clearer separation on high-performing models, highlighting that ability captures relative item difficulty and provides finer differentiation beyond raw accuracy.
+(Left) Whole-bank ability estimates show a non-linear relationship with average score and reveal clearer separation on high-performing models, highlighting that ability captures relative item difficulty and provides finer differentiation beyond raw accuracy.
 (Right) Rank comparison indicates strong but imperfect alignment (Spearman $\rho=0.86$, Kendall $\tau=0.70$), with deviations from the diagonal reflecting cases where ability-based ranking distinguishes models more effectively than accuracy alone.
+ARC
 Figure 6: Comparison of raw average scores and whole-bank ability estimates on ARC.
-(Left) Whole-bank estimates exhibit a non-linear relationship with average scores, providing clearer separation on high-performing models by accounting for item difficulty and discrimination.
+(Left) Whole-bank ability estimates exhibit a non-linear relationship with average scores, providing clearer separation on high-performing models by accounting for item difficulty and discrimination.
 (Right) Rank comparison shows strong but not perfect alignment between the two metrics (Spearman $\rho=0.91$, Kendall $\tau=0.77$), with deviations from the diagonal highlighting cases where ability-based ranking offers more informative distinctions than raw accuracy alone.
+ARC
 Figure 7: Two models with the similar average accuracy (0.713) and (0.714) on ARC nevertheless receive very different whole-bank ability estimates.
+Difficulty levels are ordered from easiest (D1) to hardest (D10).
 Model A (mera-mix-4x7B) attains a whole-bank ability rank of 270 because its correct responses are concentrated on more difficult items.
@@
 This divergence shows how IRT-based ability estimation can distinguish models that appear identical under raw accuracy by accounting for item difficulty.
+HellaSwag
 Figure 8: Two models with the same average accuracy (0.853) on HellaSwag nevertheless receive very different whole-bank ability estimates.
+Difficulty levels are ordered from easiest (D1) to hardest (D10).
 Model A (supermario_v1) attains a whole-bank ability rank of 347 because its correct responses are concentrated on more difficult items.